A Bibliography of Publications in IEEE Computational Science and Engineering

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 March 2018 Version 1.17

Title word cross-reference

3 [Ano95l, Bra97b, JYR^{+}96, ZF95]. $6 [Ano95k].

/NII [Voi94a].

1 [JW95]. 100 [Str94b]. 11th [FJSP95]. 14th [Sme98]. 16th [Leh95]. 1995 [FJSP95]. 1996 [Ric96a, Vol96]. 1997 [Bai97].

2.0 [BKMC96]. 2000 [Sto94]. 21st [Cyb96b, SGS^{+}96]. 2nd [Mil94].

3.0 [Bra96b, Bra97a].

4 [Bra96b].

5 [Bra97c, de 96].

77 [BKMC96].

90 [Bra97c, MM98b]. ’93 [Dra94e]. ’94 [Ano95i, Dra94b, Dra94c]. ’95 [Dra96a]. ’96 [Dra96b, Joh96, Pry97]. ’97 [PC98]. ’98 [Kue98].

A. [Bou97]. Abstract [Smi96a]. Academia [SGS^{+}96]. Academic [Ric94a]. Accelerate [CS98]. Accelerating [GHRW98]. Access [Ano95n, Ano95o, Voi94b]. Accommodation [Smi96d]. Accuracy [DY98]. ACES [Vol96]. ACM [FJSP95, Ano96d]. across [Mur96]. Adam [Ano95k]. Adaptation [ACMC96]. Adds [Smi95a]. Adifor [BKMC96]. Advances [Han95]. Advancing [CRS96]. Advertiser [Ano94d]. Advertiser/Product [Ano94d]. Advice [AGG^{+}97]. Affair [CMMP95]. Affect [ZF95]. After [BS97b, Ano95m, Ano96f]. Age [Fas94]. Agent [MVY97]. Ahead [Jon96, Pad96, SS96, Wij96, Zim96]. aid [Ano94k]. Air [ZDG94]. al [Han95]. Alamos [Ano95j]. Algebra [Don96]. algorithm [Ano96f]. Algorithms [Ano95h, Bai97, FAJL^{+}96, Sul95, ZKJ^{+}98, Ano95l]. Alloy [War95]. Alphabet [Voi94a]. Also [Ano96b, Ehr94d]. America [Smi95a]. Analysis [Akr98, Ano95h, Ber96b, BCM94, FKW98, Hig98, Mar98b, Moh96, Ano95l, Bou97, Coo95, Leh95, MK96]. Announcing [Ric96a]. Annual [Ano94e, Ano95e, Ano95i, Ano96c, Ano97a, Ano98b, FJSP95]. Answers [Dra94d]. Antennas [FJSP95]. Anxiety [Smi98a]. API [DM98]. APPARC [JW94]. Applications [Edw97, EH96, Kah94a, MVZ98, SB96, Taf95, Wri95, Ano94k, Koe98]. Applied [O’L97]. Applying [CD98]. Approach [CM95]. Approximation [PFT^{+}95]. Archaeology [Smi95b]. Architecture [LG94]. Architectures [Ano95h, GBS^{+}98]. archive [Ano95j]. area [Ano95m]. aren’t [Ano95k]. Army [Ano95m, Tez94]. ARPA [Dra94a]. Arrays [BDM^{+}97]. Art [Kal98, Mil94, MK96]. Articles [Ano94b, Ano94c, Ano95d, Ano95a, Ano95b, Ano95c, Ano96a, Ano96b, Cyb97a, Mey94, Sam94a, Sam94b]. ASCI [Gus98]. Asia [Kah94a, Kah95]. Aspects [CMMP95]. Associate [Sul94, Sul95, Sul97]. Association [Ano95i]. Asynchronous [LG96]. Atmospheric [FAJL^{+}96, vWC97]. atom [Ano95l]. Atomistic [NBC^{+}98]. Atoms [ZF95]. Audience [Ano96e]. Auspicious [Sam95a]. Austin [Ode95]. Automated [JWR^{+}96]. Automatic [BKMC96, vWC97]. automating [Kah94a]. Award [Ano95j, Str94b]. awarded [Ano95m]. Awards [Ano95l, Mar96, Mar98a, PM95, Str94b].

B [Wee96]. B. [Ple97, Xu97]. Back

[Sam95c]. backs [Ano95j]. Bad [HHS96]. Bahder [Wee96]. barrier [Ano95l]. Based [ABC^{+}97, ML97, MVY97, Ano94k, Bra97b]. Be [Sul97]. before [Ano96f]. Beginning [Sam95a]. Behavior [SAW97]. Behavioral [Edw97]. Belgian [CV97]. Benchmarking [EH96]. Better [Str94b]. Between [MSH94]. Biennial [Leh95]. Big [Con97, Str94b]. Billion [Ano95l]. Billion-atom [Ano95l]. Biological [BN94]. Biology [MSH94, PFT^{+}95]. Biomolecules [BKS^{+}94]. Björstad [Xu97]. blood [Ano94k]. Blue [Voi94c]. Blue-Ribbon [Voi94c]. Board [Ano94i, Ano94l, Sul97]. Bobsledding [Ehr94d]. Book [Akr98, Bra94, Cha96a, Coo95, Cyb98a, Ehr94b, Gal94, Gal97b, Gal97a, Gur97, Han95, Hig98, Hou96, Lan95, Mar98c, Mil94, Mor96, Phi96, Ple97, Saa95, Sai96, Sch95a, Van95, Vra97, Wee96, Xu97]. Books [Ano98f]. Bookshelf [Gal97b, Gal97a]. Borrowing [PC98]. Bothering [Wij96]. BRDF [ZKJ^{+}98]. Breaking [Bor97]. Breakthrough [Ano95j]. breast [Ano95l, Str94b]. Brian [Mil94]. Bridge [Ano94k]. Brinch [Van95]. broken [Ano95l]. bubble [Ano95m]. Building [Hou96, Jal94].

C [Cyb95]. CAD [Bar97]. calculated [Ano95m]. Calculus [Cyb97b]. Calendar [Ano94f, Ano94g, Ano94h, Ano95f, Ano95g, Ano98e, Ano98c, Ano98d, Ehr94c, Ehr95]. Calendar/Call [Ano95f, Ano95g, Ehr94c]. Call [Ano98e, Ehr95]. CAM [Ode95]. Can [BS96c]. Canadian [PM98]. cancer [Ano95l, Str94b]. Capabilities [Lar98]. Capability [Cyb95]. Capsid [LJT^{+}98]. Carefully [Ehr94a]. Carlo [BS97b, Mas98, DF94]. cars [Str94b]. Case [CCL^{+}98, MM98b]. Catalysis [Tru98]. Catching [Hag94, HHS96]. Causal [MS96].

Cautionary [Sul94]. CD [Ano94k]. CD-ROMs [Ano94k]. Cells [HHS96]. Cellular [STG^{+}96]. Center [Hus95, Noo95, Sim97, Ano95k, Cor94, Tay94, Tez94, Wes96]. Centers [Ber96b, Smi97b, Ano95l]. Century [Cyb96b, SGS^{+}96]. ceramics [Ano95m]. CERN [Ano95k]. Chains [Phi96]. Challenge [Cyb96b, Cyb98d]. Challenges [BG97, Gur97, Pee97, Han95]. Change [BHMF96]. Changes [Ano95l, Sul97]. chaos [Str94c]. Charles [Sai96]. Charts [Ehr94d]. Check [Rig94a]. Chemistry [DF94, FJSP95, GF95, Tru94b]. Chemkin [Bra97a]. Chief [Cyb96a, Cyb96b, CW96b, Cyb96c, Cyb97c, Cyb97d, Cyb97e, Cyb98b, Cyb98d, CH98, Cyb98c, Sam94c, Sam95a, Sam95b, Sam95c, Sul94, Sul95, Sul97]. chip [Ano95m, Bra97b]. Christoph [Hig98]. Chromodynamics [Mey94, Wil94]. Circle [Ano95d, Sul97]. Claim [Ewa96]. Climate [Kah95, Ano95m, Str94b]. Close [Smi98c]. Clustering [GCC97]. Clusters [BCM94, Gus97]. Code [vWC97, SC98]. Codes [Ell95b]. Collaborative [Bra98a]. Collisions [WPM95]. Combining [GF95]. Combustion [Ano95i]. Coming [Ehr94a, Ano95l]. Command [LG96]. comments [Str94b]. commercializes [Ano95k]. Commercializing [Pee97]. Committee [Rig94b]. Common [Bra98a]. Community [Ric97a]. companion [Ano96b]. Comparing [Pap98]. Competition [Kah94b]. Compiler [Pad96]. compilers [Bra97b]. Complementarity [FJSP95]. Complex [MS96]. CompMed [Dra94b]. Component [GBS^{+}98]. Compugotcha [Smi97a]. Computation [Ano95h, Ano97b, BGC^{+}95, BG97, BHP^{+}94, Dra94b, EPWW94, FGK^{+}95, Hig98, MM98a, Taf95, WPM95, Ano95l, Lan95]. Computational [Ano94e, Ano94j, Ano95e, Ano95h, Ano96c, Ano97a, Ano98b, ASFT94, BLC96, BDM^{+}97, CRS96, CMMP95, CM95, Cyb96e, EKP98, FJSP95, GHR94, GS97, Gar97, GVR94, Gou94, Gus98, HGBR97, JA94, Joh95, Lan96, ML97, Mar98a, Noo95, O’L97, PJB97, QAF^{+}96, Ric94a, Ric94b, Ric95, Ric96a, Ric97a, Sai96, SAW97, SHR98, Smi98d, Tru94b, USO94, Van95, VK95, Ano94k, Bra97b, Str94b, Dra94b, Dra94c, Tru97a]. Computations [Ano95h]. Computer [Ano96d, Edw97, EC96, GHR94, HD98, Iss96, JG96, Kue98, Lan96, McQ96, MSH94, Ric96b, Ric97a, Wil96a, Wil96b, Str94b, Str94c, Ano96b]. Computers [SS96]. Computing [Ano95i, BS96a, BS96b, BS96c, BS97b, BS97a, BS98, Bra97c, Cyb96b, CW96b, Cyb97b, CH98, Ehr94d, Gur97, JG96, Jon96, Kah94b, Kal95, LJT^{+}98, MWHR97, Mil94, Mor96, O’L94, Pat96, PFT^{+}95, Ric95, Saa95, Sch95a, Smi96a, Smi97c, Tez94, Voi94a, Zim96, Bra94, Leh95, Tru97b]. Concurrent [MM98a, Saa95]. Conductance [ZF95]. Conference [Dra94a, FJSP95, Kue98, Tru97b, Ano94k, Ano95h, Ano96d, BG94, Vol96]. Conferences [Ano95h, Ano95i, Ano96d, Ano97b, Bai97, BD96, Con97, Dra94d, Dra94b, Dra94c, Dra94e, Dra96a, Dra96b, Edw97, FJSP95, GP94, Gus97, Joh96, KE98, Koe98, Kue98, Leh95, Mas98, MK96, Pry97, PC98, Ric96a, Ric96b, SB96, Sme98, Tru94a, Tru97a, Tru97b, Tru98, Vol96, de 96, Koe98]. Confidence [Cyb95, Ell95a, Jal94]. Configuration [FHST98]. Connections [BS96b]. Considerations [VK95]. Consolidation [Smi96a]. Consulting [AGG^{+}97]. Contact [MSH94]. Content [GS97]. Context [JG96]. Control [BDM^{+}97, LG96, Bra97b]. Convection [JYR^{+}96]. Converges [Ric97a]. Converging [Dra94d]. Cooperate [Mur96]. cooperation [Str94b]. Cornell [Cor94]. Corporate [JiI95]. corrected [Ano96f]. Course [Akr98]. Cover [Dos95, Ehr94a, GJGLB95, Hag94, HHS96, PS95, SB94, Woo94, Zhu95]. Crack [Abr97]. Crash [FKW98]. Cray [Ber96a, Ewa96, Smi95a, Sul97]. Cray/SGI [Smi95a]. Creating [Lar98]. Critics [Ano95d]. CRS4 [QAF^{+}96]. Crunching [Peh95]. Cryo [LJT^{+}98]. Cryo-Electron [LJT^{+}98]. CS [Ric96a, Ste97]. CS&E [Sam95b]. CSE [AGG^{+}97, Ano96e, Apt97, Bor97, BG94, CV97, Cyb94, Cyb97d, Cyb97e, Cyb98d, Cyb98c, Dru98, Ell95a, Ell95b, GS97, Gar97, Gro96, Hus95, JW94, Jal94, JW95, Kah94a, Kah94b, Kah95, Kal98, MMP95, Mar96, Mar98a, Min98, MM98b, MS96, Ode95, Pan95, PM95, Peh95, PM98, Ric96b, Ric97a, Rig94b, Sam95b, San96, SGS^{+}96, Sch95b, Smi96b, Sol96, Ste97, USO94, Voi94a, Voi94b, Voi94c, Wah96, Zac98]. CSEP [USO94]. CSIRO [Dra94b]. Cup [Smi95a]. Curved [ZM98]. Cycle [Zhu95].

D [Ano95l, Bra97b, JYR^{+}96, ZF95]. DAGS’94 [Dra94c]. damage [Ano95m]. Danny [Ano97c]. Data [Ano95h, Apt97, BS96a, MM98a, PS95, SN94, Wij96, Bra97b]. database [Ehr94d]. David [Cyb96d, Gur97, Mor96]. debris [Ehr94d]. Decentralized [LG96]. Decision [MVY97]. Decomposition [Xu97]. Decoupled [Gué95]. Deep [ZF95]. Deep-Submicron [ZF95]. Deer [ABC^{+}97]. Defect [Ehr94d]. Defense [Gou94]. Deficit [Smi98a]. Deformable [Gué95]. Demystified [Mor96]. Design [BDM^{+}97, DO96, FGK^{+}95, Kue98, LPH96, Tru97a]. Designing [Hou96]. Desktop [OGP97]. Detailed [BN94]. Details [Con97]. detect [Ano95l, Str94b]. Detecting [CM95]. Detection [ZDD97]. Determination [MM98a]. Determining [LJT^{+}98]. developers [Str94b]. Developing [BRRM95, GBS^{+}98, Mar97]. Development [FHST98]. Devices [SAW97, ZF95]. DEVS [ZMKB97]. Diego [Tay94]. difference [Str94b]. Different [PFT^{+}95]. Differential [AKR^{+}97, Akr98, Ano95h, Vra97]. Differentiation [BKMC96]. director [Ano95k]. Discover [Ano94k]. Discrete [Bag98, Cyb97b]. Discrete-Event [Bag98]. disk [Ano95m]. Distributed [Ano95m, BG97, Cla97, Cyb98d, GBS^{+}98, Gus97, HDS^{+}95, LG94, ZDD97]. Distributed- [Cla97]. Distributed-Memory [HDS^{+}95]. DNA [Ano94k]. Do [Cyb96d, McQ96, Str94c]. Docking [LFH^{+}97]. DoD [Ano95k]. DoE [Mar96]. Does [War95]. Doing [Cyb97d, Str94a]. Domain [Xu97]. Double [Ano95m]. Dowd [Bra94]. Down [Ano97b, Cla98a, Ehr94a, Smi98b, Str94b]. Downsizing [Moh96]. Drama [Smi95a]. drives [Ano95m]. Drug [Tru97a]. Drug-Design [Tru97a]. drugs [Str94b]. Durations [BKS^{+}94]. Dynamic [Cyb97b, Cyb98a]. Dynamical [Bou97]. Dynamics [Abr97, Dra94b, HDS^{+}95, LPH96, QAF^{+}96, SB96].

e-mails [Bra96b]. Earth [Kin95, Leh95]. earthquakes [Ehr94d]. Ecological [FRN94]. Ecology [ABC^{+}97, BCM94, Ehr94d]. Econoflu [Smi98a]. Ecosystems [KKLK97]. Edge [ZDD97]. edited [Han95]. Edition [Mil94]. Editor [Cyb94, Ric94b, Cyb96a, Cyb96b, CW96b, Cyb96c, Cyb96e, Cyb97c, Cyb97d, Cyb97e, Cyb98b, Cyb98d, CH98, Cyb98c, EKP98, GW98, Joh95, Kal95, Mok97, O’L94, PRSE95, Sam94c, Sam95a, Sam95b, Sam95c, Str94a, Sul94, Sul95, Sul97, Web97, Wri95]. Editor-in-Chief [Cyb96a, Cyb96b, CW96b, Cyb96c, Cyb97c, Cyb97d, Cyb97e, Cyb98b, Cyb98d, CH98, Cyb98c, Sam94c, Sam95a, Sam95b, Sam95c, Sul94, Sul95, Sul97]. Editorial [Ano94i, Ano94l, Sul97]. Editors [CRS96, HGBR97, Rig94a, RW97, Tru94b]. Education [AGG^{+}97, BG94, CV97, Cyb97e, Gar97, MMP95, Mar96, ML97, Mar98a, Mey94, Min98, Pan95, PM95, PM98, Rig94b, SGS^{+}96, Wah96, Wil94, Zac98]. Educational [BG94, Ode95, Rig94b, USO94]. Educators [Mar96]. Effects [FAJL^{+}96, GF95]. Efficient [KL98, Sul95, Str94c]. Efforts [PM98]. Einstein [Hag94]. elected [Ehr94d]. Electrical [CI95]. Electromagnetic [Dru98, KL98]. Electromagnetics [CM95, GVR94, Str94c, Taf95, VK95, Bra97b]. Electron [LJT^{+}98, WPM95]. Electron-Molecule [WPM95]. electronic [Ano95j]. Electrostatic [KL98]. Element [Coo95, Fas94, Ano95k]. embedded [Bra97b]. embedded-processor [Bra97b]. Emerging [O’L94, Taf95]. Emil [Bra98a]. Emphasizing [Kah94a]. Emulating [Cla97]. Enabled [CD98]. End [Cyb97c]. Engine [Zhu95]. Engineer [Sol96]. Engineering [Cyb96e, CH98, Dou96, EC96, EKP98, Gar97, JA94, ML97, Ric94a, Ric94b, Ric96a, Ric97a, Str94b, Web97, Ano98b, Ano94e, Ano94j, Ano95e, Ano96c, Ano97a]. Engineers [Iss96, JG96, Mar98c, McQ96, Wil96a, Wil96b, Ano94k, Wee96]. Eniac [EC96]. enough [Ano95k]. Entertainment [Ano97c]. Environment [AKR^{+}97, Ano95i, ZMKB97, Ano95j]. Environmental [BRRM95, Leh95]. Environments [GHR94, HGBR97, JWR^{+}96, RB96]. EPCC [Min98]. Equation [GHRW98]. Equations [AKR^{+}97, DY98, Vra97, Akr98, Ano95h]. Eric [Saa95]. Errors [Hat97]. Establishes [Ode95]. Estimating [EPWW94]. Euphoritopia [Smi98a]. Europe [ASFT94, Ell95a, Ell95b, JW94, Jal94, JW95, Sch95b]. Europort [Ell95b, Jal94, JW95]. Europort-1 [JW95]. Europort2 [Ell95a]. Evaluation [Sam95c]. Evaluations [CS98]. Eve [Ano95k]. Event [Bag98]. Everglades [ABC^{+}97]. Examples [Ano96f, ACMC96, PFT^{+}95]. Excellence [PM95]. Exchange [BDG^{+}95]. Experience [EPWW94, Wah96]. Experimental [CW96b]. Experiments [Hat97]. Exploration [Ano95h]. Exploring [SN94, TCD^{+}96]. Explosive [Zhu95]. Extra [Ano96e, Ano96f]. eyes [Str94b].

F [Saa95]. facility [Ano95l]. Fare [Sam95b]. Farewell [Sam95b]. farming [Str94b]. Fast [Ano95l, CS98, DY98, FAJL^{+}96, GHRW98, GW98, GS98, Ano96f, SC98]. FD [Taf95]. FD-TD [Taf95]. Feasibility [Gus98]. Feedback [PRSE95]. Field [War95, Ano95m]. Fields [PC98]. film [Ano95m]. Finance [BLC96, Smi98d]. Find [BS96c]. Finding [Bra98a, Str94c]. Fingerprint [GCC97]. Fingerprints [HHS96]. Finite [Fas94, Ano95k, Coo95]. Finite-Element [Fas94, Ano95k]. First [Akr98, Min98, Ric96a, Ric97a, Sto94, Ano95i, FJSP95, Ric96b]. five [Ano95k]. five-year [Ano95k]. Flannery [Mil94]. Flows [Woo94]. Fluid [Dra94b, QAF^{+}96]. FMC [Str94c]. Focus [Cla98a, Cla98b, Voi94a, Voi94b, Voi94c]. Forecast [vWC97]. forget [Iss96]. Fortran [Bra97a, Mar98c, BKMC96, Bra97c, MM98b]. Forum [Str94b]. Forward [Sam95c]. Foster [Hou96]. Four [Pan95, Zhu95]. Four-Stroke [Zhu95]. Fracture [Abr97]. Free [Bra97c]. Freeze [War95]. Frequencies [GHRW98]. Frontiers [Pry97]. Fujitsu [Ano95i]. full [Ric95]. Future [ASFT94, Don96, Ehr94d, Gur97, Ric95, Ric97b, Smi97c]. Fuzzy [ACMC96, JWR^{+}96].

G [Mey94]. Gate [Ano94k]. Geller

[Mar98a]. GEMS [BRRM95]. Generation [Cla98b, vWC97, Ano95j, Ano95h]. Generators [Bra98b]. Genome [Ehr94d]. Geographic [Bra98b]. Geomagnetic [Ano95m]. Geometric [Bar97, GPA97, RW97, WR97]. Geometry [Ano95h, FJSP95]. Geophysical [PS95]. Get [Cyb96d, Wah96, Str94b]. Gets [BG94, Cla98a, Ano94k]. Gibbs [Lan95]. Glaciology [Fas94]. Global [ZKJ^{+}98, Str94b]. Globally [Kal95]. Goes [Sol96, Ste97]. gold [Str94c]. Golden [Ano94k]. Good [Cyb96d, Str94c]. Gory [Con97]. grants [Ano95k]. graphics [Bra97b]. Grid [Ano95h]. Grids [Woo94]. Gropp [Xu97]. Ground [Bra98a]. Groundwater [EPWW94]. Group [EH96]. Grow [PM98]. growing [Ehr94d]. Guest [HGBR97, EKP98, GW98, RW97]. Guide [Ano94j, Ano96b]. Guided [BGC^{+}95]. guitars [Ano95m]. Gulf [Ano95j]. Guys [HHS96].

H [Mil94]. Hand [Smi98c]. Hands [Wah96]. Hands-On [Wah96]. Hansen [Van95]. Hardly [Cyb97c]. Hardware [Smi96a, Smi97c]. Hashing [Bar97, GPA97, RW97, WR97]. Haus [Ano95m]. Heap [BS96a]. Heating [Ano95i]. Helmholtz [GHRW98, GS98]. help [Ano94k]. Herman [Ano95m]. Hierarchical [Smi95b]. High [BG94, CCL^{+}98, Cyb95, EH96, Jon96, Kah94b, SKP96, Smi96a, Smi96c, Tez94, Voi94a, ZMKB97, ZKJ^{+}98, Zim96, Bra97b, Bra94, Edw97, FJSP95, Gur97, Mas98, Mor96]. High-Level [CCL^{+}98]. High-Performance [EH96, Jon96, Kah94b, SKP96, Smi96a, Smi96c, Voi94a, ZMKB97, ZKJ^{+}98, Zim96, Bra97b, Bra94, Edw97, FJSP95, Mas98]. Highway [Voi94b]. Hillis [Ano97c]. Hirlam [CW96a]. hold [Ano94k]. honor [Str94b]. honored [Ano94k, Str94c]. Honors [KE98]. HotGAMS [Bra96b]. HPC [Ber96b, Ehr94d, Ell95a, Jal94, Kah94a, McM97, Smi96b]. HPCC [Dra94a, Ano95k, BDG^{+}95, Voi94a]. HPCC/IITA [Voi94a]. HPF [Str94b]. Humanity [Smi97a]. Humphries [Bou97].

Ian [Hou96]. ICASE [Hus95]. Ice [Fas94]. ICIAM [Leh95]. ICIP [PC98]. Idaho [Str94b]. IEEE [Ano94e, Ano94j, Ano95e, Ano96c, Ano97a, FJSP95, Ric96b, Sme98, Ano98b, Ric96a, Ric97a, Ste97]. IEEE-CS [Ric96a]. IES [KL98]. If [Sam94b]. II [Ano95h]. IITA [Voi94b]. ILAY [BD96]. Illinois [SC98]. Image [Ano95l, Cyb94, GPA97]. Image-analysis [Ano95l]. imagery [Ano96f]. Images [FAJL^{+}96]. Imaging [BGC^{+}95, CI95, Cru94, Ano95m, Ple97]. impacts [Str94b]. Impedance [CI95]. Implementations [ZDD97]. improve [Ano95m]. Improving [FHST98]. Impurity [ZF95]. Increasing [Ell95a]. Index [Ano94e, Ano94d, Ano95e, Ano96c, Ano97a, Ano98b]. Indiana [Ric96a]. Individual [ABC^{+}97]. Individual-Based [ABC^{+}97]. Indoor [FGK^{+}95]. Industrial [Apt97, EH96, Ell95a, Str94c]. Industry [Apt97, Bor97, DM98, Dru98, Mar98a, San96, SGS^{+}96, Kah94a]. Industry-Standard [DM98]. Informant [Bra96b]. Information [Bra98b, Voi94b]. Infosphere [Cha96b]. Infrastructure [Mar97, Smi97b]. Initiative [MMP95]. Innovation [PM95]. Intelligence [Cyb98d]. Intelligent [MVY97]. Interactive [Ano95h, CRS96, JYR^{+}96, KKLK97, LFH^{+}97, OGP97, STG^{+}96, TCD^{+}96]. Interconnection [SS96]. Interested [Rig94b]. Interface [Leh95, Ano95k]. Interfaces [Smi95a, Smi95b, Smi96b, Smi96a, Smi96c, Smi96d, Smi97b, Smi97a, Smi97c, Smi98a, Smi98c, Smi98b, Smi98d]. Interior [Kin95]. Interleave [BS98]. International [Ano95i, FJSP95, LG94, Kue98, Tru97b]. Internet [Ano95k]. Interview [Ano97c, Cyb95, Cyb96d]. Int’l [Ano95h]. Introducing [Ano94l, CH98]. Introduction [CRS96, Cyb96e, EKP98, Gra95, GW98, HGBR97, Joh95, RW97, Sai96, USO94, Zac98, Mar98c, Phi96]. Inverse [Joh95, Kah94a]. Issues [CI95, ZDD97]. Italy [QAF^{+}96]. Iterative [BD96, Cha96a, KE98]. IWCE [GP94].

J [Cyb96d, Gur97, Han95, Phi96]. J. [Vra97]. Japan [JiI95, Kah94b, Smi96b]. Java [Bra97a]. Job [Pap98]. Job-Management [Pap98]. John [Cyb95]. Journey [Kin95]. Jr. [KE98]. Jupiter [Str94b].

Kalos [Ehr94d]. Kaufmann [Wee96]. Kenneth [Mey94]. Kevin [Bra94]. Kiwis [Smi95a]. Knocking [Cyb96a]. Knowledge [Cyb98d]. Krylov [Dru98]. Kuck [Cyb96d, Gur97].

l [Ano95m]. Lab [Str94b]. Laboratories [Ano95p]. Laboratory [OGP97]. labs [Str94c]. Lag [TCD^{+}96]. Land [BHMF96]. Land-Use [BHMF96]. Landscape [BCM94]. Langley [Hus95]. Language [MVZ98]. Languages [Bag98, BHP^{+}94, EKP98, Kal98, Zim96, Sme98]. laptops [Bra98b]. Large [Cyb96b]. Large-Scope [Cyb96b]. Larger [BKS^{+}94]. Larry [Mar98a]. Lawrence [Han95]. layer [Ano95m]. leech [Ano95m]. let [Smi98b]. Letter [Mok97, Web97]. Letters [Ano95d, Ewa96, PRSE95]. Leukemia [CM95]. Level [CCL^{+}98]. Levels [PFT^{+}95]. Libraries [Don96, RB96]. LIfe [Dra94b, Str94b]. Line [Str94c, USO94]. Linear [Cha96a, Don96]. Linear-Algebra [Don96]. linked [Ric95]. Lion [Mar98b]. Literature [Ano95n, Ano95o]. Loan [Sai96]. Lockheed [Str94b]. Long [ZDG94]. Long-Range [ZDG94]. Longer [BKS^{+}94]. Looking [Jon96, Pad96, Sam95c, SS96, SN94, Wij96, Zim96, Ano95m]. Lookup [BS96c]. Loshin [Mor96]. Low [GHRW98, Str94b]. Low-life [Str94b]. lubricants [Ano95m]. Lucas [BHMF96].

M [Bou97, KE98, Vra97]. M. [Ple97]. Mac [Ano94k]. Mac-based [Ano94k]. machine [Ano94k]. Machines [HDS^{+}95]. MacNeal [Ano94k]. MacNeal-Schwendler [Ano94k]. MacSciTech [Ano95i]. magazine [Ano94k]. magnetism [Ano95m]. mails [Bra96b]. Make [BS97a]. makes [Str94b]. Making [BS96b, Str94c]. Management [FHST98, Pap98]. Managing [Str94a]. Mania [Cyb98c]. manipulation [Bra97b]. Mantle [JYR^{+}96]. manufacturing [Ano94k, Ano95j, Ano95h]. Many [Tru94b]. Maple [Bra96b]. Market [Sol96]. Markov [Phi96]. Marks [BG94]. Mass [Mar97]. Mass-Produce [Mar97]. Match [BS97a]. Matching [GCC97]. Matchup [Smi95a]. Materials [Ano95h, NBC^{+}98]. Math [AGG^{+}97, Ano97b]. Matheij [Vra97]. Mathematica [Bra96b, Iss96, Wee96]. Mathematical [Bor97]. Mathematics [O’L97, Sai96]. Mathworks [Bra97b]. Matlab [Bra97c, Bra97a, Web97]. Maxwell [DY98]. may [Ano94k, Ano95m]. Me [BS97a]. Mean [Smi97b]. Meaning [PS95]. Means [Gus97]. Mechanics [Abr97, GF95, Gou94, Dra94c]. Medal [Ano95m]. Media [Bou97]. Medical [GPA97, Wri95]. Medicine [Joh95]. Meditation [Sul97]. Meeting [Ano95i, FJSP95]. Meetingplace [Cyb96c]. Meets [HD98]. Members [Ano94l]. Memorium [Ber96a]. Memory [Cla97, DM98, HDS^{+}95]. MEMS [BG97, BDM^{+}97, Mar97, Pee97, SAW97]. Merger [Cyb98c]. Metal [Bor97, War95]. Method [BGC^{+}95, CS98, Fas94, GS98]. Methods [BLC96, BDM^{+}97, Cha96a, DY98, DF94, Dou96, GVR94, GHRW98, GW98, Hig98, KE98, SAW97, VK95, BD96]. Micromanipulator [BDM^{+}97]. Microprocessor [Pat96]. Microscopy [LJT^{+}98]. Midsummer [Smi95b]. Military [Jon96, LG96]. Millennium [Smi98c]. million [Ano94k, Ano95k]. millisecond [Str94c]. Mining [Apt97]. Model [FHST98, War95, Str94b]. modeled [Ano95m]. Modeling [ABC^{+}97, Ano95j, Ano95m, BHMF96, BKS^{+}94, Bor97, BRRM95, Coo95, EPWW94, Fas94, FRN94, STG^{+}96, ZMKB97, ZDG94, vWC97]. Models [Bag98, PS95, ZKJ^{+}98]. Modern [Lan95]. Modernizing [Jon96]. Modular [FRN94]. Molecular [Abr97, GF95, GJGLB95, HDS^{+}95, LPH96, LFH^{+}97, MSH94, PFT^{+}95, SB96]. Molecule [WPM95]. Molecules [Tru94a]. Molenaar [Vra97]. Monitoring [Ano95h, MS96]. Monte [BS97b, Mas98, DF94]. Move [McM97, Tru94a]. Moving [Smi96d]. MRI [BGC^{+}95]. MRI-Guided [BGC^{+}95]. MTA [FKW98]. Multidisciplinary [MVZ98]. Multigrid [Dou96, SB96]. Multilevel [GS98]. Multimedia [Cyb97e]. multiple [Iss96]. Multipole [CS98, DY98, GHRW98, GW98, GS98]. Multiscale [SN94].

Nanostructured [NBC^{+}98]. Nanotechnology [Cla98c]. NASA [Ano95k, Dra94a, Hus95]. National [Ano95p, BDG^{+}95, Gou94, JiI95, Str94b, Ano95m]. Natural [Ano95i]. navigation [Kah94a]. Needs [SGS^{+}96, SAW97, Sul95]. NEOS [CMM98]. NERSC [Sim97]. Net [ACMC96, MWHR97]. nets [Str94b]. NetSolve [CD98]. Network [BN94, CD98, Gro96]. Network-Enabled [CD98]. Networks [Cyb96e, DO96, Gus97, MS96, SS96, de 96, Ano95h]. Neural [ACMC96, Cyb96e, DO96, Str94b, de 96]. Neuro [JWR^{+}96, Cyb98a]. Neuro-Dynamic [Cyb98a]. Neuro-Fuzzy [JWR^{+}96]. Neuronal [BN94]. New-Order [Smi97b]. News [Akr98, Ano94k, Ano95j, Ano95k, Ano95l, Ano95m, Ano97d, Ber96b, Ber94a, Ber94b, Ber94c, Ber95a, Ber95b, Ber95c, Ber95d, Ber96c, Bou97, Bra94, Bra96a, Bra96b, Bra97c, Bra97b, Bra97a, Bra98b, Bra98a, Bri96, Cha96a, Coo95, Cyb98a, Ehr94b, Ehr94d, Ehr94e, Gal94, Gal97b, Gal97a, Gur97, Han95, Hig98, Hou96, Lan95, Mar98b, Mar98c, McM97, Mil94, Moh96, Mor96, Pap98, Phi96, Ple97, Saa95, Sai96, Sch95a, Str94b, Str94c, Van95, Vra97, Wee96, Xu97]. Next [Ano95j, Cla98b, Ano95k]. Next-generation [Ano95j]. Night [Smi95b]. Nightmare [Smi95b]. Niño [Smi98b]. NIST [Bra96b]. Nos [Ano94e, Ano95e, Ano96c, Ano97a]. Note [Sul94]. NSF [Ano95j, Ano95k, Ano95l, Ber96b, Ehr94d, Smi97b, Voi94c]. NT [Bra97b, Gus97]. NT-based [Bra97b]. Number [Bra98b]. Numbers [Peh95]. Numerical [Akr98, Bou97, Hig98, Kin95, Leh95, MK96, Phi96, Leh95, Mil94].

Object [Ano95i]. Object-Oriented

[Ano95i]. Objects [Bar97]. Ocean [Mur96]. October [Ric96a]. ODE [Bra97a]. off [ZM98]. offer [Ano94k]. offers [Ano95k]. Oil [Dru98, EPWW94, Ano95m]. On-Line [Str94c, USO94]. One [GS97]. OpenMP [Bra97a, DM98]. opens [Ano95k]. Opportunity [Cyb96a]. Optical [BGC^{+}95, HHS96]. Optimization [ACMC96, FGK^{+}95, Wri95]. Optimized [Ano95j]. Option [BLC96]. Order [Smi97b]. Ordinary [Vra97]. Organic [GF95]. Oriented [Ano95i]. Origin [Bri96]. Other [PC98]. Our [Cyb97f]. Outline [Pad96]. Overview [WR97].

P [Mil94]. P. [Xu97]. Packages [Bra97c]. Painful [Smi96a]. Panel [Ehr94d, Voi94c]. Papers [Ano95f, Ano95g, Ehr94c, Ehr95]. Paradigm [Gar97]. Paradigms [Van95]. Parallel [ABC^{+}97, Abr97, Ano95h, Ano95i, Bag98, BCM94, Bra98b, BN94, CCL^{+}98, Cyb96d, Ell95a, Hou96, Jal94, JiI95, JYR^{+}96, LJT^{+}98, Leh95, O’L94, RC94, SS96, STG^{+}96, Str94b, Tru97b, Van95, WPM95, Zim96, Ano94k, Leh95, Sch95a]. Parallelism [Pan96]. Parallelized [Ell95b]. Parallelizing [HDS^{+}95]. Parameter [ACMC96, GCC97]. Parameters [EPWW94]. Part [Iss96, JG96, Jon96, McQ96, Pad96, SS96, Wij96, Wil96b, Zim96]. Partial [AKR^{+}97, BS96c]. particles [Ano94k]. partners [Str94b]. Partnership [Smi97b]. Passing [Sul97]. Past [ASFT94]. PathForward [Cla98a]. Payments [LG94]. PDEs [JWR^{+}96]. peace [BS98]. Pellpack [MWHR97]. Call [Ano95f, Ano95g, Ehr94c]. Doer [GHR94]. IITA [Voi94a]. NII [Voi94a]. Product [Ano94d]. SGI [Smi95a]. Pentium [Bra97b, CMMP95]. Performance [Cyb95, Cyb96d, EH96, Gur97, Jon96, Kah94b, Mor96, Rig94a, SKP96, Smi96a, Smi96c, Tez94, Voi94a, ZMKB97, ZKJ^{+}98, Zim96, Bra97b, Bra94, Edw97, FJSP95, Mas98]. Permpar [Sch95b]. Perspective [Apt97, Kal95, O’L94, Pap98, Wri95]. Perspectives [Cyb94, EC96, Ric94b, Rig94a, Tru94b]. Petaflops [Bai97, Dra94a]. Pharmaceutical [Mar98a]. Phase [War95]. Phase-Field [War95]. Physical [Iss96, JG96, McQ96, Wil96a, Wil96b]. Physicists [de 96]. Physics [Lan95]. Picture [Con97]. Pieces [BS98]. Pittsburgh [Wes96]. plan [Ano95k]. planes [Str94c]. Plans [Voi94c]. Playing [Ano97c]. Point [MSH94]. Pointing [Cla98c]. Pollutants [ZDG94]. port [Ano94k]. Portable [Bra98b]. Portait [Abr97]. power [Bra98b]. Practical [FGK^{+}95]. practicality [Ano94k]. Practice [Vra97]. Practices [Leh95]. Prediction [DO96, Ehr94d]. Prescription [GS98, Mar98a]. Prescriptions [BS96a, BS96b, BS96c, BS97b, BS97a, BS98, Cyb97b]. Present [ASFT94]. Press [Mil94]. Pricing [BLC96]. Principles [McC96, Leh95]. private [Str94b]. prize [Ano94k, Ehr94d]. Problem [AKR^{+}97, GHR94, GBS^{+}98, HGBR97, JWR^{+}96, KKLK97, RB96, Kah94a]. Problem-Solving [AKR^{+}97, GHR94, HGBR97, JWR^{+}96, KKLK97, RB96, Kah94a]. Problems [Joh95, San96, FJSP95]. Processing [Ano95h, Cyb94, JiI95, LG94, Ano95i]. processor [Bra97b]. Produce [Mar97]. Product [Ano97d, Ber94a, Ber94b, Ber94c, Ber95a, Ber95b, Ber95c, Ber95d, Ber96c, Bra96a, Ehr94e, GS97]. Products [Bra97c, Bra97a]. Program [Gus98, Ode95]. Programme [Min98]. Programming [BHP^{+}94, CCL^{+}98, Cla97, Cyb97b, DM98, Gro96, Kal98, Leh95, McC96, Van95, Wij96, Zac98, Cyb98a, Leh95]. Programs [BKMC96, BG94, HDS^{+}95, Ode95, Ric94a, Rig94b, USO94, Hou96]. Progress [Smi96a]. Project [CW96a, CW96a, JW94, Sch95b]. Promising [McM97]. Propagation [FJSP95]. Properties [SC98]. Proposals [Ber96b]. Pros [Bra97b]. Prospects [BGC^{+}95]. Protein [Ano95m, LPH96, Ano95l]. PSE [MWHR97]. PSEs [Bra97b]. Public [Str94b]. Public-private [Str94b]. pulsars [Str94c]. Puppy [Gar97]. Purdue [Ric96a, Ste97]. PVM [Str94b]. PVR [SKP96].

Quantum [GF95, Mey94, Wil94, Tru98]. Quick [SHR98]. Quick-Turnaround [SHR98]. Quickly [BS96c, Ano95m].

R [Bou97, Lan95]. R&D

[Ano95j, Ano95k, Kah95, Str94b]. R&D-100 [Str94b]. R. [Vra97]. Radio [Cru94]. Ramp [Voi94b]. Random [Ano95n, Ano95o, Bra98b]. Randomly [SB94]. Range [ZDG94]. Rapid [Abr97, DO96]. Reader [Ano94m]. Real [Cla98c, EH96, LG94, San96]. Real-World [Cla98c]. Reality [Rig94a, SN94, TCD^{+}96, Ano94k]. Really [Smi97b]. Recipes [Mil94]. Reciprocity [Smi96c]. Recognize [PM95]. Reconnaissance [GJGLB95]. record [Bra97b]. Recursive [Dos95]. Redesigning [Mey94, Wil94]. Referees [Ano94a]. Reflection [Sto94, ZM98]. Registration [GPA97]. Reinventing [Sim97, Taf95]. Rejection [Smi96d]. Release [Bra96b]. Reliability [Bor97]. Removing [FAJL^{+}96]. Rendering [SKP96]. Repair [Bar97]. Reply [Wil96b]. Report [Ano95p, CW96a, Cor94, Hus95, Noo95, QAF^{+}96, Sim97, Tay94, Tez94, Wes96]. Representation [Gué95]. Requirements [SC98]. Rescuing [Smi97a]. Research [Apt97, Bra98a, Ric95, Tez94, Wah96, Ano94k, Ano95j, Ehr94d, Str94c, Ewa96]. Resolution [Str94b]. Resource [Gou94]. Response [Iss96, JG96, Lan96, McC96, McQ96, Voi94c, Web97]. responses [Ano96e]. Results [Ell95b, Zhu95]. Retrieval [ZKJ^{+}98]. reversals [Ano95m]. Review [Hig98]. Reviewers [Ano97e, Ano98a, Cyb97f]. Reviews [Akr98, Bou97, Bra94, Bra96b, Bra97b, Bra97a, Bra98b, Bra98a, Bri96, Cha96a, Coo95, Cyb98a, Gal97b, Gal97a, Gur97, Han95, Hig98, Hou96, Lan95, Mar98c, McM97, Mil94, Mor96, Pap98, Phi96, Ple97, Saa95, Sai96, Sch95a, Van95, Vra97, Wee96, Xu97]. Rex [Smi95b]. Ribbon [Voi94c]. riblets [Str94c]. Rightsizing [Moh96]. Roadmap [Pad96]. Roads [Tru94b]. Rocket [HD98]. Rocketry [HD98]. Roggemann [Ple97]. ROMs [Ano94k]. Rosenblum [Han95]. Roundtable [Ano96e]. Run [AGG^{+}97, Str94b].

Saad [Cha96a]. safer [Str94c]. Salute [Ano97e, Cyb97f]. San [Tay94]. Sandia [Ano94k, Ano95p]. Sarpa [Bra98a]. Satellite [FAJL^{+}96, Ano96f]. Saul [Mil94]. scalability [Ano95j]. Scalable [FRN94, Kal95, Bra98b]. ScaLapack [Bra97b]. Scales [BKS^{+}94]. Scaling [SC98]. scans [Bra96b]. SCCM [AGG^{+}97]. Scholarship [Min98]. Schwendler [Ano94k]. Science [Ano94e, Ano94j, Ano95e, Ano96c, Ano96d, Ano97a, Ano98b, Cyb96e, Cyb97c, CH98, Dou96, EC96, EKP98, GHR94, GS97, Gar97, HD98, HGBR97, JA94, Lan96, ML97, Mar98a, MSH94, O’L97, Ric94a, Ric94b, Ric95, Ric96a, Ric97a, Sai96, USO94, Van95, Web97, Ano95m, Ano95h, Ano95m]. Sciences [Ano95i, Dra94b, Edw97, Leh95]. Scientific [BHP^{+}94, CW96b, Ehr94d, EPWW94, GBS^{+}98, Hat97, Koe98, MWHR97, Mil94, Pat96, RB96, Ric97b, Saa95, Zac98, Ano95h, Bra97a, Han95]. Scientist [Cha96b]. Scientists [Iss96, JG96, Lan96, McQ96, Wee96, Wil96a, Wil96b, Ano94k, Str94b, Mar98c]. SciNapse [AKR^{+}97]. SciTools [Joh96]. Scope [Cyb96b]. SDSC [Ano95l, Wah96]. Searching [RC94]. seeks [Ano95k]. Selective [GVR94]. Sequoia [Sto94]. Series [Bri96, DO96]. Server [CD98, CMM98, MWHR97, Ano95k]. Seventh [Ano95h]. Seymour [Ber96a, Sul97]. SGI [Bri96]. Shapes [Smi97c, Ano95l]. Shaping [Peh95]. Shared [Cla97, DM98]. Shared-Memory [DM98]. Should [Iss96, JG96, Lan96, McQ96, Wil96a, Wil96b]. SIAM [Ano95h, FJSP95, Dra94c]. Side [Ano97c]. Signal [Cyb94]. Similar [RC94]. Simplify [Cla97]. simulated [Ano95m]. Simulating [Ano94k, GF95, LG96, SAW97, Zhu95]. Simulation [Bag98, BN94, KL98, KKLK97, Lar98, Leh95, MVY97, NBC^{+}98, SHR98, STG^{+}96, ZMKB97, ZF95, Ano95m, Kah94a]. Simulation-Based [MVY97]. Simulations [TCD^{+}96]. Single [Ano95m]. Single-layer [Ano95m]. Site [Ano95p, Cor94, Hus95, Noo95, QAF^{+}96, Sim97, Tay94, Tez94, Wes96]. smallest [Ano95m]. Smart [Ano95m, PC98]. SMD [LPH96]. Smith [Xu97]. Smithsonian [Str94b]. Social [Leh95]. Society [FJSP95, Ric96b, Ric97a, Str94c]. Software [Bra97c, BDG^{+}95, Hat97, Hig98, PJB97, RB96, Ric97b, Web97, Bra97a, Str94b, Leh95]. Solidification [War95]. Solution [JWR^{+}96, Phi96]. Solvent [GF95]. Solver [SC98, Ano94k]. Solving [AKR^{+}97, GHR94, GBS^{+}98, HGBR97, JWR^{+}96, KKLK97, RB96, Kah94a]. Some [ACMC96]. Soup [Voi94a]. Space [Ehr94d]. Sparse [Cha96a]. SPEC [EH96]. Speedup [Dos95, Ehr94d]. Spherical [MM98a]. Spherical-Virus [MM98a]. Spirit [Smi96b]. Spline [CS98]. Splines [Gué95]. Spotlight [Cla98c]. Springer [Hig98]. Springer-Verlag [Hig98]. SPRNG [Bra98b]. Stalk [LFH^{+}97]. Standard [DM98, Bra97a]. Standards [CW96b]. Stands [Ewa96]. Stanford [AGG^{+}97]. stars [Ano95j]. State [Kal98, MK96]. Steel [Peh95]. Steering [BHP^{+}94, CRS96, LPH96, PJB97]. Step [JWR^{+}96]. Stephan [Wee96]. Stewart [Phi96]. Stiff [Kah94b]. Stock [Jon96, Pad96, SS96, Smi98c, Wij96, Zim96]. Strategies [LJT^{+}98, PJB97]. Stream [Ano95j]. Street [Sol96]. Stress [Coo95]. Strings [RC94]. Stroke [Zhu95]. Structural [Gou94]. Structure [LJT^{+}98, MM98a]. Structures [ASFT94, Noo95, Wij96, Ano95m]. Stuart [Bou97]. Student [AGG^{+}97]. Student-Run [AGG^{+}97]. Students [Wah96]. Studies [Van95]. Study [MM98b]. studying [Ano95m]. Stuff [BS96c]. Submicron [ZF95]. Submissions [Ano94j]. Subspaces [Dru98]. substitute [Ano94k]. Summer [Min98]. Sums [BS96c]. Sun [McM97]. super [Ehr94d]. Supercomputer [Ano95m, Sim97, Str94b, Tay94]. supercomputers [Ano95k].

Supercomputing [Cla98b, Wes96, Ano95k, Dra94c, Dra94e, Dra96a, Dra96b]. Superfine [Woo94]. Supernetworking [Ehr94d]. Support [EC96, JWR^{+}96, MVY97, MVZ98]. Surface [Gué95, ZM98]. Survey [Ano94m, EPWW94, GVR94]. Symbolic [Bra97c]. Symposium [Ano95h, FJSP95, KE98, Sme98, Dra94a, Dra94b, FJSP95]. Synapse [SB94]. System [BHMF96, BRRM95, FRN94, LFH^{+}97, Bra97b]. Systems [Bou97, Bra98b, FGK^{+}95, MS96, Pap98, PJB97, Ric97b, Rig94a, Cha96a, Gur97].

T [Hou96, Mil94, Hat97]. Taking

[Jon96, Pad96, SS96, Wij96, Zim96]. TD [Taf95]. Teach [Iss96, JG96, Lan96, McC96, McQ96, Web97, Wil96a, Wil96b]. Teamwork [O’L97]. Teasing [PS95]. Technical [Ano97c, Bra96b, Rig94b]. Technoculture [Smi96d]. Technology [ASFT94, Bra97c, Bra97b, Bra97a, Bra98b, Bra98a, Bri96, McM97, Pad96, Pap98, Ano95i, Noo95]. telemanufacturing [Ano95l]. Tera [FKW98]. Teraflops [Smi95b, Ano95j]. Teramac [Cla98c]. Teukolsky [Mil94]. text [Ric95]. thar [Str94c]. them [Str94c]. Theme [CRS96, Cyb96e, Joh95]. Theory [Cor94, Vra97]. There [Str94c]. Thin [Ano95m]. Thin-film [Ano95m]. Thinker [GHR94]. Thinker/Doer [GHR94]. Third [Ano95i, Ano95i, GP94]. Thomas [Wee96]. Thoughts [Ste97]. Three [Sto94]. Time [BS97b, DO96, LG94, O’L94, TCD^{+}96]. Tips [McC96]. Tomography [BGC^{+}95]. Tomorrow [vWC97]. Tool [STG^{+}96, Bra97b, Wee96]. Toolbox [Bra97a]. Toole [Cyb95]. Tools [McC96, Iss96, Mas98]. Trade [Smi96c]. Tradition [Bor97]. Training [ACMC96]. Transformation [GCC97]. Transient [GVR94]. Transport [ZDG94]. Travel [Tru94b]. Tree [BS96c]. Tree-Lookup [BS96c]. Trends [JiI95, SS96]. Turbulence [Ple97]. Turbulent [Woo94]. Turmoil [Smi98d]. Turnaround [SHR98]. Turns [EC96]. Tutorial [FJSP95, SB96].

UCES [MMP95]. Ueberhuber [Hig98]. UK [Ano95k]. Unbroken [Sul97]. Undergraduate [MMP95]. Understanding [Ehr94d]. Universe [Cru94]. Universities [CV97]. University [Ric96a, JA94]. updates [Bra97b]. Upsizing [Moh96]. Use [BHMF96, Iss96]. User [Leh95, Pap98]. Users [Ano95i]. Using [Abr97, Bar97, CS98, CM95, GCC97, Gué95, GPA97]. UT [Ode95]. Utah [JA94]. UVA [Noo95].

V [Bra96b]. Variables [Gué95]. Variations [Cyb97b]. vehicle [Kah94a]. Velde [Saa95]. Verifiability [Gus98]. Verlag [Hig98]. version [Str94b]. Vetterling [Mil94]. via [MWHR97, USO94]. Viewpoint [GS97]. Viral [LJT^{+}98]. Virtual [Ano94k, HD98, LFH^{+}97, SN94, TCD^{+}96]. Virus [MM98a]. viruses [Str94b]. Vision [Sam95c]. Visual [BHP^{+}94, LPH96, OGP97, Sme98, Ano95h]. Visualization [CRS96, Han95, JYR^{+}96, MVY97, MM98a]. Visualizing [PS95, ZM98]. VLSI [Ano95m]. Volume [Ano94e, Ano95e, Ano96c, Ano97a, Ano98b, SKP96].

W [Hig98]. W. [Phi96, Xu97]. Walking [SB94]. Wall [Sol96]. Warming [Kah95]. Warp [O’L94]. Watching [Tru94a]. Wavelengths [Cru94]. Wavelets [Gra95]. Waves [Hag94]. Way [Cla98c]. Weather [vWC97]. Web [ML97, Bra96b, Ano96e, Ano96f, Bra96b, Bra97c, MWHR97]. Web-Based [ML97]. Week [McQ96]. welcomes [Str94b]. Well [Sam95b]. Welsh [Ple97]. Whither [Ric94b]. Who [Sul95].

Will [Smi97b, Sul97]. William

[Lan95, Mil94]. Wilson

[Iss96, JG96, Mey94, Web97]. Win [Mar96, Smi95a]. Wind [Mar98b]. wins [Ano94k]. Wireless [FGK^{+}95]. WISE [FGK^{+}95]. without [Wij96]. Wolfram [Bra96b, Bra97b]. Won’t [McQ96]. Work [Ano96e, Cla98a, MS96, Peh95, San96, SGS^{+}96, Sol96, Str94c]. Workbench [KKLK97]. WorkPlace [Bra97a]. Works [Sch95a]. Workshop [Ano95i, BD96, Leh95, Ric96a, Ric96b, Ric97a, Ste97, Ano95i, Bai97, FJSP95, Tru97a]. Workshops [Ano95h, Ano95i, Ano96d, Ano97b, Bai97, BD96, Con97, Dra94a, Dra94d, Dra94b, Dra94c, Dra94e, Dra96a, Dra96b, Edw97, FJSP95, GP94, Gus97, Joh96, KE98, Koe98, Kue98, Leh95, Mas98, MK96, Pry97, PC98, Ric96a, Ric96b, SB96, Sme98, Tru94a, Tru97a, Tru97b, Tru98, Vol96, de 96, Dra94a]. workstation [Bra97b]. World [Cla98c, Bra97b, PC98]. Wrongsizing [Moh96].

Year [Pat96, Ano95k]. Years

[Sto94, Min98]. Yield [Ell95b]. yields [Ano95m]. Young [KE98]. Yousef [Cha96a].

Zeroes [Smi98c]. ZPL [CCL^{+}98].

References

Abbott:1997:PIB

[ABC^{+}97] Catherine A. Abbott, Michael W. Berry, E. Jane Comiskey, Louis J. Gross, and Hank-Kwang Luh. Parallel individual-based modeling of everglades deer ecology. IEEE Computational Science & Engineering, 4(4):60–72, October/December 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1997/pdf/c4060.
pdf;   http://www.computer.
org/cse/cs1998/c4060abs.htm.

Abraham:1997:PCR

[Abr97]  Farid F. Abraham.  Por-

tait of a crack: Rapid fracture mechanics using parallel molecular dynamics. IEEE Computational Science & Engineering, 4(2):66–77, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2066.pdf;   http:
//www.computer.org/cse/cs1998/
c2066abs.htm.

Arabshahi:1996:FPA

[ACMC96] Payman Arabshahi, Jai J. Choi, Robert J. Marks, II, and Thomas P. Caudell. Fuzzy parameter adaptation in optimization: Some neural net training examples. IEEE Computational Science & Engineering, 3(1):57–65, Spring 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c1057abs.htm.

Aczon:1997:CES

[AGG^{+}97] Melissa Aczon, Martin Gander, Margot Gerritsen, Tony Shardlow, and Ronnie Sircar. CSE education: SCCM advice: Studentrun math consulting at Stanford. IEEE Computational Science & Engineering, 4(1):7–9, January/ March 1997. CODEN ISCEE4.

ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c1007.pdf.

Akers:1997:SPS

[AKR^{+}97] Robert L. Akers, Elaine Kant, Curtis J. Randall, Stanly Steinberg, and Robert L. Young. Sci-Napse: a problem-solving environment for partial differential equations. IEEE Computational Science & Engineering, 4(3):32–42, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c3032.pdf;   http:
                                        //www.computer.org/cse/cs1998/
                                        c3032abs.htm.

Akrivis:1998:BNR

[Akr98] Georgios Akrivis. Book news & reviews: A First Course in the Numerical Analysis of Differential Equations. IEEE Computational Science & Engineering, 5(2):103, April/June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c2101.pdf.

Anonymous:1994:R

[Ano94a] Anonymous. 1994 referees. IEEE Computational Science & Engineering, 1(4):96, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:AAa

[Ano94b] Anonymous. About the articles. IEEE Computational Science & Engineering, 1(1):10, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:AAb

[Ano94c] Anonymous. About the articles. IEEE Computational Science & Engineering, 1(4):14–15, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:API

[Ano94d] Anonymous. Advertiser/product index. IEEE Computational Science & Engineering, 1(1):96, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:AII

[Ano94e] Anonymous. Annual index 1994:

IEEE Computational Science & Engineering, volume 1, nos. 1– 4. IEEE Computational Science & Engineering, 1(4):94–95, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:Ca

[Ano94f] Anonymous. Calendar. IEEE Computational Science & Engineering, 1(1):92–93, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:Cb

[Ano94g] Anonymous. Calendar. IEEE Computational Science & Engineering, 1(3):95, Fall 1994. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:Cc

[Ano94h] Anonymous. Calendar. IEEE Computational Science & Engineering, 1(4):91–92, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:EB

[Ano94i] Anonymous. The editorial board. IEEE Computational Science & Engineering, 1(1):4–7, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:ICS

[Ano94j] Anonymous. IEEE Computational Science & Engineering guide for submissions. IEEE Computational Science & Engineering, 1(1):94–95, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:NMS

[Ano94k] Anonymous. In the news: MacNeal-Schwendler to port applications to parallel machine; A substitute for blood?; simulating 400 million particles; virtual reality may aid manufacturing; Sandia’s parallel solver wins practicality prize; Golden Gate Bridge gets help from computational engineers; computational DNA research honored by Discover magazine; Mac-based scientists hold conference, offer CD-ROMs. IEEE Computational Science & Engineering, 1(4):87–90, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:INE

[Ano94l] Anonymous. Introducing new editorial board members. … IEEE Computational Science & Engineering, 1(2):93, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1994:RS

[Ano94m] Anonymous. Reader survey.

IEEE Computational Science & Engineering, 1(3):87, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:AAa

[Ano95a] Anonymous. About the articles. IEEE Computational Science & Engineering, 2(2):16–17, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:AAb

[Ano95b] Anonymous. About the articles. IEEE Computational Science & Engineering, 2(3):10–11, Fall 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:AAc

[Ano95c] Anonymous. About the articles. IEEE Computational Science & Engineering, 2(4):22–23, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:AAL

[Ano95d] Anonymous. About the articles; letters; critics circle. IEEE Computational Science & Engineering, 2(1):15–16, Spring 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:AII

[Ano95e] Anonymous. Annual index 1995: IEEE Computational Science & Engineering, volume 2, nos. 1– 4. IEEE Computational Science & Engineering, 2(4):94–95, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:CCPa

[Ano95f] Anonymous. Calendar/call for papers. IEEE Computational Science & Engineering, 2(1):95, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:CCPb

[Ano95g] Anonymous. Calendar/call for papers. IEEE Computational Science & Engineering, 2(2):95, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:CWM

[Ano95h] Anonymous. Conferences & workshops: Manufacturing and Computational Geometry; Visual Data Exploration and Analysis II; Scientific Computation and Differential Equations; Visual Monitoring of Computations — Interactive Grid Generation; Int’l Symposium on Parallel Architectures and Algorithms and Networks; Seventh SIAM Parallel Processing Conference; Int’l Symposium on Materials Science. IEEE Computational Science & Engineering, 2(2):82–89, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:CWO

[Ano95i] Anonymous. Conferences & workshops: Object-Oriented Computing in the Natural Sciences ’94; MacSciTech Users Association Third Annual Meeting; Combustion and Environment and Heating Technology; Third Fujitsu Parallel Computing Workshop; First International Workshop on Parallel Processing. IEEE Computational Science & Engineering, 2(1):90–93, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:NMG

[Ano95j] Anonymous. In the news: Modeling the gulf stream NSF backs los alamos electronic archive; R&D stars; breakthrough award; from teraflops to scalability; nextgeneration research environment; optimized manufacturing and more. IEEE Computational Science & Engineering, 2(2):74–77, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:NNG

[Ano95k] Anonymous. In the news: NSF grants $6 million; DoD’s five-year R&D plan; HPCC seeks new director; CERN offers supercomputing server; UK opens supercomputing center; NASA commercializes finite-element interface; Adam on the Internet; next year: Eve; when supercomputers aren’t enough. IEEE Computational Science & Engineering, 2 (1):85–87, 89, Spring 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:NST

[Ano95l] Anonymous. In the news: SDSC telemanufacturing facility; fast computation of 3D protein shapes; NSF centers: Changes coming?; image-analysis algorithms detect breast cancer; awards; billion-atom barrier broken; more. IEEE Computational Science & Engineering, 2(3):77– 79, Fall 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:NTF

[Ano95m] Anonymous. In the news: Thin-film lubricants may damage disk drives; protein structures calculated quickly; supercomputer looking for oil; modeling ceramics may improve yields; VLSI chip modeled after a leech; US Army studying imaging science; geomagnetic field reversals simulated; single-layer magnetism; National Medal of Science awarded to Herman A. Haus; distributed climate simulation; double bubble area is the smallest;l smart guitars. IEEE Computational Science & Engineering, 2 (4):82–84, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:LRAa

[Ano95n] Anonymous. Literature: Random access. IEEE Computational Science & Engineering, 2 (2):93, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:LRAb

[Ano95o] Anonymous. Literature: Random access. IEEE Computational Science & Engineering, 2 (4):89, Winter 1995. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1995:SRS

[Ano95p] Anonymous. Site report: Sandia National Laboratories. IEEE Computational Science & Engineering, 2(2):10–15, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1996:AA

[Ano96a] Anonymous. About the articles. IEEE Computational Science & Engineering, 3(3):17, Fall 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1996:AAG

[Ano96b] Anonymous. About the articles: Also: a guide to companion articles in Computer. IEEE Computational Science & Engineering, 3 (1):16–17, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1996:AI

[Ano96c] Anonymous. Annual index 1996: IEEE Computational Science & Engineering, volume 3, nos. 1– 4. IEEE Computational Science & Engineering, 3(4):80–82, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1996:CWA

[Ano96d] Anonymous. Conferences & workshops: ACM Computer Science Conference. IEEE Computational Science & Engineering, 3(2):91–92, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Anonymous:1996:WEA

[Ano96e] Anonymous. Web extra: Audience responses to “CSE at work” roundtable. IEEE Computational Science & Engineering, 3(2):??, Summer 1996. CODEN ISCEE4.

ISSN 1070-9924 (print), 1558-190X (electronic). URL extras/

c2004x.htm.

Anonymous:1996:WEE

[Ano96f] Anonymous. Web extra: Examples of before and after satellite imagery corrected by the fast algorithm. IEEE Computational Science & Engineering, 3(2):??, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL extras/

c2066x.htm.

Anonymous:1997:AI

[Ano97a] Anonymous. Annual index 1997: IEEE Computational Science & Engineering, volume 4, nos. 1–4. IEEE Computational Science & Engineering, 4(4):94–96, October/December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c4094.pdf.

Anonymous:1997:CWM

[Ano97b] Anonymous. Conferences & workshops: Math and computation down under. IEEE Computational Science & Engineering, 4(2):84–85, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2082.pdf.

Anonymous:1997:IDH

[Ano97c] Anonymous. Interview: Danny Hillis: Playing with the technical side of entertainment.

IEEE Computational Science & Engineering, 4(4):6–8, October/December 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://dlib.computer.
                                        org/cs/books/cs1997/pdf/c4006.
                                        pdf.

Anonymous:1997:PN

[Ano97d] Anonymous. Product news. IEEE Computational Science & Engineering, 4(2):79–81, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c2078.pdf.

Anonymous:1997:SR

[Ano97e] Anonymous. Salute to 1997 reviewers. IEEE Computational Science & Engineering, 4(4):97, October/December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c4097.pdf.

Anonymous:1998:R

[Ano98a] Anonymous. 1998 reviewers. IEEE Computational Science & Engineering, 5(4):3, October/December 1998. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://dlib.computer.
                                        org/cs/books/cs1998/pdf/c4003.
                                        pdf.

Anonymous:1998:AII

[Ano98b] Anonymous. Annual index: IEEE Computational Science and Engineering, volume 5. IEEE Computational Science & Engineering, 5(4):83–87, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c4083.pdf.

Anonymous:1998:Ca

[Ano98c] Anonymous. Calendar. IEEE Computational Science & Engineering, 5(3):95, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c3095.pdf.

Anonymous:1998:Cb

[Ano98d] Anonymous. Calendar. IEEE Computational Science & Engineering, 5(4):82, October/ December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1998/pdf/c4082.
pdf.

Anonymous:1998:CC

[Ano98e] Anonymous. Call & calendar. IEEE Computational Science & Engineering, 5(1):86– 87, January/March 1998. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1998/pdf/c1086.
pdf.

Anonymous:1998:NB

[Ano98f] Anonymous. New books. IEEE Computational Science & Engineering, 5(1):83, January/March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c1083.pdf.

Apte:1997:CID

[Apt97] Chidanand Apté. CSE in industry: Data mining: An industrial research perspective. IEEE Computational Science & Engineering, 4(2):6–9, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c2006.pdf.

Argyris:1994:CST

[ASFT94] John Argyris, Ioannis St. Doltsinis, Gerhard Frik, and Lazarus Tenek. Computational structures technology in Europe: Past, present, and future. IEEE Computational Science & Engineering, 1(1):43–54, Spring 1994. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Bagrodia:1998:PLD

[Bag98]  Rajive L. Bagrodia.  Par-

allel languages for discreteevent simulation models. IEEE Computational Science & Engineering, 5(2):27–38, April/ June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2027.pdf;   http:
//www.computer.org/cse/cs1998/
c2027abs.htm.

Bailey:1997:CWP

[Bai97] David Bailey. Conferences & workshops: 1997 Petaflops Algorithms Workshop. IEEE Computational Science & Engineering, 4(2):82–84, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2082.pdf.

Barequet:1997:UGH

[Bar97]  Gill Barequet.  Using geo-

metric hashing to repair CAD objects. IEEE Computational Science & Engineering, 4(4):22–28, October/December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c4022.pdf;   http:
//www.computer.org/cse/cs1998/
c4022abs.htm.

Berry:1994:PAC

[BCM94] Michael Berry, Jane Comiskey, and Karen Minser. Parallel analysis of clusters in landscape ecology. IEEE Computational Science & Engineering, 1(2):24–38, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Benzi:1996:CWI

[BD96] Michele Benzi and Craig Douglas. Conferences & workshops: ILAY Workshop on Iterative Methods. IEEE Computational Science & Engineering, 3 (3):84–86, Fall 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Browne:1995:NHS

[BDG^{+}95] Shirley Browne, Jack Dongarra, Stan Green, Keith Moore, Tom Rowan, Reed Wade, Geoffrey Fox, Ken Hawick, Ken Kennedy, Jim Pool, Rick Stevens, Bob Olson, and Terry Disz. The national HPCC software exchange. IEEE Computational Science & Engineering, 2(2): 62–69, Summer 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c2062abs.htm.

Bohringer:1997:CMD

[BDM^{+}97] Karl F. Böhringer, Bruce R. Donald, Noel C. MacDonald, Gregory T. A. Kovacs, and John W. Suh. Computational methods for design and control of MEMS micromanipulator arrays. IEEE Computational Science & Engineering, 4(1):17–29, January/ March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1017.pdf;   http:
//www.computer.org/cse/cs1998/
c1017abs.htm.

Berry:1994:PNa

[Ber94a] Michael Berry. Product news. IEEE Computational Science & Engineering, 1(2):89–91, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Berry:1994:PNb

[Ber94b] Michael Berry. Product news.

IEEE Computational Science & Engineering, 1(3):82–84, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Berry:1994:PNc

[Ber94c] Michael Berry. Product news.

IEEE Computational Science & Engineering, 1(4):81–83, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Berry:1995:PNa

[Ber95a] Michael Berry. Product news.

IEEE Computational Science & Engineering, 2(1):82–84, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Berry:1995:PNb

[Ber95b] Michael Berry. Product news.

IEEE Computational Science & Engineering, 2(2):78–81, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Berry:1995:PNc

[Ber95c] Michael Berry. Product news.

IEEE Computational Science & Engineering, 2(3):80–83, Fall 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Berry:1995:PNd

[Ber95d] Michael Berry. Product news. IEEE Computational Science & Engineering, 2(4):78–81, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Berlin:1996:MSC

[Ber96a] F. Brett Berlin. In memorium: Seymour Cray, 1925–1996. IEEE Computational Science & Engineering, 3(4):90, 92, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Berlin:1996:NAN

[Ber96b] F. Brett Berlin. News analysis: NSF HPC centers: Proposals are in! IEEE Computational Science & Engineering, 3 (3):75, Fall 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Berry:1996:PN

[Ber96c] Michael Berry. Product news. IEEE Computational Science & Engineering, 3(1):92–93, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Bramley:1994:EPC

[BG94] Randall Bramley and Efstratios Gallopoulos. Educational programs in CSE: CSE Education Conference gets high marks. IEEE Computational Science & Engineering, 1(2):76–79, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Berlin:1997:DMN

[BG97] Andrew A. Berlin and Kaigham J.

Gabriel. Distributed MEMS: New challenges for computation. IEEE Computational Science & Engineering, 4(1):12–16, January/ March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1012.pdf;   http:
//www.computer.org/cse/cs1998/
c1012abs.htm.

Barbour:1995:MGO

[BGC^{+}95] Randall L. Barbour, Harry L.

Graber, Jenghwa Chang, San-Lian S. Barbour, Ping C. Koo, and Raphael Aronson. MRI-guided optical tomography: Prospects and computation for a new imaging method. IEEE Computational Science & Engineering, 2(4): 63–77, Winter 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c4063abs.htm.

Berry:1996:LSM

[BHMF96] Michael W. Berry, Brett C.

Hazen, Rhonda L. MacIntyre, and Richard O. Flamm. Lucas: a system for modeling landuse change. IEEE Computational Science & Engineering, 3 (1):24–35, Spring 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c1024abs.htm.

Burnett:1994:TVP

[BHP^{+}94] Margaret Burnett, Richard Hossli, Tim Pulliam, Brian VanVoorst, and Xiaoyang Yang. Toward visual programming languages for steering scientific computation. IEEE Computational Science & Engineering, 1(4):44–62, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Bischof:1996:AAD

[BKMC96] Christian Bischof, Peyvand Khademi, Andrew Mauer, and Alan Carle. Adifor 2.0: Automatic differentiation of Fortran 77 programs. IEEE Computational Science & Engineering, 3(3):18–32, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c3018abs.htm.

Board:1994:MBL

[BKS^{+}94] John A. Board, Jr., Laxmikant V. Kalé, Klaus Schulten, Robert D. Skeel, and Tamar Schlick. Modeling biomolecules: Larger scales, longer durations. IEEE Computational Science & Engineering, 1 (4):19–30, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Barucci:1996:CMF

[BLC96] Emilio Barucci, Leonardo Landi, and Umberto Cherubini. Computational methods in finance: Option pricing. IEEE Computational Science & Engineering, 3(1):66–80, Spring 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c1066abs.htm.

Brettle:1994:DPS

[BN94] Dean Brettle and Ernst Niebur.

Detailed parallel simulation of a biological neuronal network. IEEE Computational Science & Engineering, 1(4):31–43, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Borucki:1997:CIB

[Bor97] Len Borucki. CSE in industry: Breaking with tradition in mathematical metal reliability modeling. IEEE Computational Science & Engineering, 4(3):10–12, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c3010.pdf.

Bountis:1997:MNR

[Bou97] Anastasios Bountis. Media news & reviews: Dynamical Systems and Numerical Analysis by A. M. Stuart and A. R. Humphries. IEEE Computational Science & Engineering, 4(2):86–87, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c2086.pdf.

Bramley:1994:BNR

[Bra94] Randall Bramley. Book news & reviews; High-Performance Computing, by Kevin Dowd. IEEE Computational Science & Engineering, 1(4):84–85, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Bramley:1996:PN

[Bra96a] Randall Bramley. Product news. IEEE Computational Science & Engineering, 3(2):96–98, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Bramley:1996:TNR

[Bra96b] Randall Bramley. Technical news & reviews: Maple V Release 4; Mathematica 3.0 from Wolfram; HotGAMS from NIST; Informant scans Web and e-mails you; more on the web. … IEEE Computational Science & Engineering, 3(3):80, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). Discusses a Java-based front end for the Guide to Available Mathematical Software, with a pointer

                                        to URL http://math.nist.gov/
                                        HotGAMS/.

Bramley:1997:TNRb

[Bra97a] Randall Bramley. Technology news & reviews: Chemkin software; OpenMP Fortran Standard; ODE toolbox for Matlab; Java products; Scientific WorkPlace 3.0. IEEE Computational Science & Engineering, 4(4):75–78, October/December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c4075.pdf.

Bramley:1997:TNRa

[Bra97b] Randall Bramley. Technology news & reviews: New embedded-processor chip; control system PSEs from Wolfram, Mathworks; new data manipulation tool; ScaLapack updates; NT-based 3D graphics workstation; high-performance compilers for Pentium Pros; world record in computational electromagnetics. IEEE Computational Science & Engineering, 4(3):88–90, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c3088.pdf.

Bramley:1997:TNF

[Bra97c] Randall Bramley. Technology news: Fortran 90 news; free software; symbolic computing packages; Matlab 5; Web products. IEEE Computational Science & Engineering, 4(1):87, 89–90, January/ March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1087.pdf;   http:

                                        //www.cs.berkeley.edu/~xiaoye;
                                         http://www.cs.cornell.edu/
                                        home/vavasis/qmg-home.html;
                                         http://www.globus.org/sage;
                                        http://www.macsyma.com/;
                                        http://www.mathworks.com/;
                                        http://www.netlib.org/benchmark/
                                        linpackjava;   http://www.
                                        netlib.org/scalapack;   http:
                                        //www.ucmp.berkeley.edu/subway/
                                        phylogen.html;   http://www.
                                        vni.com/products/wpd/jnl/jnl_
                                        1_0.html;   http://www.wolfram.
                                        com/look/cse.

Bramley:1998:TNRb

[Bra98a] Randall Bramley. Technology news & reviews: Emil Sarpa on collaborative research: Finding common ground. IEEE Computational Science & Engineering, 5(3):8–10, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c3008.pdf.

Bramley:1998:TNRa

[Bra98b] Randall Bramley. Technology news & reviews: Geographic information systems; Scalable Parallel Random Number Generators — SPRNG; portable power for laptops. IEEE Computational Science & Engineering, 5(1):13–15, January/March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c1013.pdf.

Brimley:1996:TNR

[Bri96] Randall Brimley. Technology news & reviews: The SGI Origin series. IEEE Computational Science & Engineering, 3(4):93– 94, Winter 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Bruegge:1995:DGE

[BRRM95] Bernd Bruegge, Erik Riedel, Armistead G. Russell, and Gregory J. McRae. Developing GEMS: An environmental modeling system. IEEE Computational Science & Engineering, 2(3):55–68, Fall 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3055abs.htm.

Beichl:1996:CPH

[BS96a] Isabel Beichl and Francis Sullivan. Computing prescriptions: a heap of data. IEEE Computational Science & Engineering, 3 (2):11–15, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Beichl:1996:CPM

[BS96b] Isabel Beichl and Francis Sullivan. Computing prescriptions: Making connections. IEEE Computational Science & Engineering, 3(3):9–12, Fall 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Beichl:1996:CPT

[BS96c] Isabel Beichl and Francis Sullivan. Computing prescriptions:

Tree-lookup for partial sums or: How can I find this stuff quickly? IEEE Computational Science & Engineering, 3(1):13–15, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Beichl:1997:CPMb

[BS97a] Isabel Beichl and Francis Sullivan. Computing prescriptions: Make me a match. IEEE Computational Science & Engineering, 4(4):88–93, October/December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c4088.pdf.

Beichl:1997:CPMa

[BS97b] Isabel Beichl and Francis Sullivan. Computing prescriptions: (monte carlo) time after time. IEEE Computational Science & Engineering, 4(3):91–94, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c3091.pdf.

Beichl:1998:CPI

[BS98] Isabel Beichl and Francis Sullivan. Computing prescriptions: Interleave in peace, or interleave in pieces. IEEE Computational Science & Engineering, 5(2):92–96, April/ June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2092.pdf.

Chamberlain:1998:CHL

[CCL^{+}98] Bradford L. Chamberlain, Sung-Eun Choi, Christopher Lewis, Lawrence Snyder, W. Derrick Weathersby, and Calvin Lin. The case for high-level parallel programming in ZPL. IEEE Computational Science & Engineering, 5(3):76–86, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c3076.pdf;   http:
//www.computer.org/cse/cs1998/
c3076abs.htm.

Casanova:1998:ANN

[CD98] Henri Casanova and Jack Dongarra. Applying NetSolve’s network-enabled server. IEEE Computational Science & Engineering, 5(3):57–67, July/ September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1998/pdf/c3057.
pdf;   http://www.computer.
org/cse/cs1998/c3057abs.htm.

Cybenko:1998:ECI

[CH98] George Cybenko and Lewis M.

Holmes. From the Editorin-Chief: Introducing computing in science & engineering. IEEE Computational Science & Engineering, 5(4):1–2, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c4001.pdf. This

editorial announces the merger of IEEE Computational Science and Engineering with the American Institute of Physics journal Computers in Physics, effective January 1999. The new journal will be known as Computing in Science and Engineering.

Chan:1996:BNR

[Cha96a] Tony F. Chan. Book news & reviews: Iterative Methods for Sparse Linear Systems by Yousef Saad. IEEE Computational Science & Engineering, 3(4):87–90, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Chandy:1996:SI

[Cha96b] K. Mani Chandy. The scientist’s infosphere. IEEE Computational Science & Engineering, 3(2):43– 44, Summer 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Cheney:1995:IEI

[CI95] Margaret Cheney and David Isaacson. Issues in electrical impedance imaging. IEEE Computational Science & Engineering, 2(4):53–62, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (elec-

                                        tronic).  URL http://www.
                                        computer.org/cse/cs1998/c4053abs.
                                        htm.

Clarke:1997:ESM

[Cla97]  Jerry A. Clarke.  Emulat-

ing shared memory to simplify distributed- memory programming. IEEE Computational Science & Engineering, 4(1):55–62, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1055.pdf;   http:
//www.computer.org/cse/cs1998/
c1055abs.htm.

Clark:1998:FPG

[Cla98a] David Clark. Focus: PathForward gets down to work. IEEE Computational Science & Engineering, 5(1):84–85, January/ March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c1084.pdf.

Clark:1998:FSN

[Cla98b] David Clark. Focus: Supercomputing: The next generation. IEEE Computational Science & Engineering, 5(4):79–81, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c4079.pdf.

Clark:1998:STP

[Cla98c] David Clark. Spotlight: Teramac:

Pointing the way to real-world nanotechnology. IEEE Computational Science & Engineering, 5(3):88–90, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c3088.pdf.

Colton:1995:NAD

[CM95] David Colton and Peter Monk. A new approach to detecting leukemia: Using computational electromagnetics. IEEE Computational Science & Engineering, 2(4):46–52, Winter 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c4046abs.htm.

Czyzyk:1998:NS

[CMM98] Joseph Czyzyk, Michael P. Mesnier, and Jorge J. Moré. The NEOS server. IEEE Computational Science & Engineering, 5(3):68–75, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c3068.pdf;   http:
//www.computer.org/cse/cs1998/
c3068abs.htm.

Coe:1995:CAP

[CMMP95] Tim Coe, Terje Mathisen, Cleve Moler, and Vaughan Pratt. Computational aspects of the Pentium affair. IEEE Computational Science & Engineering, 2(1):18–30, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http:/

/ieeexplore.ieee.org/stamp/

stamp.jsp?tp=&arnumber=372929;
http://www.computer.org/cse/
cs1998/c1018abs.htm.

Conroy:1997:CWB

[Con97] John M. Conroy. Conferences & workshops: From the big picture to the gory details. IEEE Computational Science & Engineering, 4(3):86–87, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c3086.pdf. Review

of the SIAM Annual Meeting in July 1997.

Cook:1995:BNR

[Coo95] Robert D. Cook. Book news & reviews: Finite Element Modeling for Stress Analysis. IEEE Computational Science & Engineering, 2(2):90–91, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Corbit:1994:SRC

[Cor94] Margaret Corbit. Site report:

The Cornell Theory Center. IEEE Computational Science & Engineering, 1(4):10–13, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Chen:1996:TEI

[CRS96] Jim X. Chen, David Rine, and Horst D. Simon. Theme Editors’ introduction: Advancing interactive visualization and computational steering. IEEE Computational Science & Engineering, 3 (4):13–17, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Crutcher:1994:IUR

[Cru94] Richard M. Crutcher. Imaging the universe at radio wavelengths. IEEE Computational Science & Engineering, 1(2):39– 49, Summer 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Chen:1998:UFM

[CS98] Fang Chen and David Suter. Using a fast multipole method to accelerate spline evaluations. IEEE Computational Science & Engineering, 5(3):24– 31, July/September 1998. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1998/pdf/c3024.
pdf;   http://www.computer.
org/cse/cs1998/c3024abs.htm.

Cuyt:1997:CEC

[CV97] Annie Cuyt and Brigitte Verdonk. CSE in education: CSE at Belgian universities. IEEE Computational Science & Engineering, 4(4):79–83, October/December 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1997/pdf/c4079.
pdf.

Cats:1996:PRH

[CW96a] Gerard Cats and Lex Wolters.

Project report: The Hirlam Project. IEEE Computational Science & Engineering, 3(4):4– 7, Winter 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Cybenko:1996:ECS

[CW96b] George Cybenko and Harry Wijshoff. From the Editor-in-Chief: Scientific standards for experimental computing. IEEE Computational Science & Engineering, 3(3):1, Fall 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Cybenko:1994:EPC

[Cyb94] George Cybenko. Editor’s perspectives: CSE for signal and image processing. IEEE Computational Science & Engineering, 1 (2):92–934, Summer 1994. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Cybenko:1995:IJC

[Cyb95]  George Cybenko.  Interview:

John C. Toole: High performance, high confidence, high capability. IEEE Computational Science & Engineering, 2(3):6– 9, Fall 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Cybenko:1996:ECK

[Cyb96a] George Cybenko. From the Editor-in-Chief: Knocking opportunity or opportunity knocking?

IEEE Computational Science & Engineering, 3(4):1, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Cybenko:1996:ECL

[Cyb96b] George Cybenko. From the Editor-in-Chief: Large-scope computing: a challenge for the 21st Century. IEEE Computational Science & Engineering, 3 (2):1, 10, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Cybenko:1996:NEC

[Cyb96c] George Cybenko. From the new Editor-in-Chief: Meetingplace. IEEE Computational Science & Engineering, 3(1):1, 95, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Cybenko:1996:IDJ

[Cyb96d] George Cybenko. Interview: David J. Kuck: What is good parallel performance, and how do we get it? IEEE Computational Science & Engineering, 3(1):81– 85, Spring 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Cybenko:1996:TEI

[Cyb96e] George Cybenko. Theme Editor’s introduction: Neural networks in computational science and engineering. IEEE Computational Science & Engineering, 3(1):36–43, Spring 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c1036abs.htm.

Cybenko:1997:AA

[Cyb97a] George Cybenko. About the articles. IEEE Computational Science & Engineering, 4(1):9, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1007.pdf.

Cybenko:1997:CPD

[Cyb97b] George Cybenko. Computing prescriptions: Dynamic programming: a discrete calculus of variations. IEEE Computational Science & Engineering, 4(1):92–97, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1092.pdf.

Cybenko:1997:ECHa

[Cyb97c] George Cybenko. From the Editor-in-Chief: Hardly the end of science. IEEE Computational Science & Engineering, 4(2):1, April/June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2001.pdf.

Cybenko:1997:ECHb

[Cyb97d] George Cybenko. From the Editor-in-Chief: How is CSE doing? IEEE Computational Science & Engineering, 4(4):2, October/December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c4002.pdf.

Cybenko:1997:ECM

[Cyb97e] George Cybenko. From the Editor-in-Chief: Multimedia in CSE education. IEEE Computational Science & Engineering, 4(3):1, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c3001.pdf.

Cybenko:1997:SOR

[Cyb97f] George Cybenko. Salute to our 1996 reviewers. IEEE Computational Science & Engineering, 4(1):91, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1091.pdf.

Cybenko:1998:BNR

[Cyb98a] George Cybenko. Book news & reviews: Neuro-Dynamic Programming. IEEE Computational Science & Engineering, 5(2):101–102, April/June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2101.pdf.

Cybenko:1998:EC

[Cyb98b] George Cybenko. From the Editor-in-Chief. IEEE Computational Science & Engineering, 5(2):1, April/June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2001.pdf.

Cybenko:1998:ECM

[Cyb98c] George Cybenko. From the Editor-in-Chief: Merger mania in CSE. IEEE Computational Science & Engineering, 5(3):1–3, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c3001.pdf.

Cybenko:1998:ECC

[Cyb98d] George Cybenko. From the Editor-in-Chief: The CSE challenge in knowledge and distributed intelligence. IEEE Computational Science & Engineering, 5(1):1, January/March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c1001.pdf.

deForcrand:1996:CWN

[de 96] Philippe de Forcrand. Conferences & workshops: Neural Networks for Physicists 5. IEEE Computational Science & Engineering, 3(1):90–91, Spring 1996.

CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Doll:1994:MCM

[DF94] Jim Doll and David L. Freeman. Monte Carlo methods in chemistry. IEEE Computational Science & Engineering, 1(1):22– 32, Spring 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Dagum:1998:OIS

[DM98] Leonardo Dagum and Ramesh Menon. OpenMP: An industrystandard API for shared-memory programming. IEEE Computational Science & Engineering, 5(1):46–55, January/March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c1046.pdf;   http:
//www.computer.org/cse/cs1998/
c1046abs.htm.

Drossu:1996:RDN

[DO96] Radu Drossu and Zoran Obradovic. Rapid design of neural networks for time series prediction. IEEE Computational Science & Engineering, 3(2): 78–89, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c2078abs.htm.

Dongarra:1996:FLA

[Don96] Jack Dongarra. Future linearalgebra libraries. IEEE Computational Science & Engineering, 3(2):38–40, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c2038abs.htm.

Dosanjh:1995:ACR

[Dos95] Sudip S. Dosanjh. About the cover: Recursive speedup. IEEE Computational Science & Engineering, 2(1):4–5, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Douglas:1996:MMS

[Dou96] Craig C. Douglas. Multigrid methods in science and engineering. IEEE Computational Science & Engineering, 3(4): 55–68, Winter 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c4055abs.htm.

Draper:1994:CWA

[Dra94a] Richard N. Draper. Conference & workshops: The ARPA HPCC Symposium and the NASA Petaflops Workshops. IEEE Computational Science & Engineering, 1(2):81–82, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Draper:1994:CWCb

[Dra94b] Richard N. Draper. Conferences & workshops: CompMed ’94, CSIRO LIfe Sciences Computation Symposium, Computational Fluid Dynamics ’94. IEEE Computational Science & Engineering, 1(3):93–95, Fall 1994. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Draper:1994:CWCc

[Dra94c] Richard N. Draper. Conferences & workshops: Computational Mechanics; SIAM; Supercomputing ’94; DAGS’94. IEEE Computational Science & Engineering, 1(4):78–79, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Draper:1994:CWCa

[Dra94d] Richard N. Draper. Conferences & workshops: Converging on answers. IEEE Computational Science & Engineering, 1 (3):92, Fall 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Draper:1994:CWS

[Dra94e] Richard N. Draper. Conferences & workshops: Supercomputing ’93. IEEE Computational Science & Engineering, 1(1):85–86, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Draper:1996:CWSa

[Dra96a] Richard N. Draper. Conferences & workshops: Supercomputing ’95. IEEE Computational Science & Engineering, 3(1):88–90, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Draper:1996:CWSb

[Dra96b] Richard N. Draper. Conferences & workshops: Supercomputing ’96. IEEE Computational Science & Engineering, 3(4):84–86, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Druskin:1998:CIK

[Dru98] Vladimir Druskin. CSE in industry: Krylov subspaces and electromagnetic oil. IEEE Computational Science & Engineering, 5(1):10–12, January/March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c1010.pdf.

Dembart:1998:AFM

[DY98] Ben Dembart and Elizabeth Yip. The accuracy of fast multipole methods for Maxwell’s equations. IEEE Computational Science & Engineering, 5(3):48–56, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c3048.pdf;   http:
//www.computer.org/cse/cs1998/
c3048abs.htm.

Eigenmann:1996:ETP

[EC96] Rudolph Eigenmann and George Cybenko. As Eniac turns 50: Perspectives on computer science support for science and engineering. IEEE Computational Science & Engineering, 3 (2):16–17, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c2016abs.htm.

Edwards:1997:CWH

[Edw97] Lynne K. Edwards. Conferences & workshops: High-Performance Computer Applications in the Behavioral Sciences. IEEE Computational Science & Engineering, 4(1):100–101, January/ March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c1098.pdf.

Eigenmann:1996:BRI

[EH96] Rudolf Eigenmann and Siamak Hassanzadeh. Benchmarking with real industrial applications: The SPEC highperformance group. IEEE Computational Science & Engineering, 3(1):18–23, Spring 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c1018abs.htm.

Ehrlich:1994:ACC

[Ehr94a] Miles Ehrlich. About the cover: Coming down carefully. IEEE Computational Science & Engineering, 1(2):4–5, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Ehrlich:1994:BN

[Ehr94b] Miles Ehrlich. Book news. IEEE Computational Science & Engineering, 1(1):91, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Ehrlich:1994:CCP

[Ehr94c] Miles Ehrlich. Calendar/call for papers. IEEE Computational Science & Engineering, 1(2):95, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Ehrlich:1994:NNP

[Ehr94d] Miles Ehrlich. In the news: NSF panel charts future of HPC; also: Supernetworking, super speedup; bobsledding Kalos elected; scientific computing prize; space debris; ecology research; understanding earthquakes; genome database growing; defect prediction. IEEE Computational Science & Engineering, 1(1):79–83, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Ehrlich:1994:PN

[Ehr94e] Miles Ehrlich. Product news.

IEEE Computational Science & Engineering, 1(1):87–90, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Ehrlich:1995:CCP

[Ehr95] Miles Ehrlich. Calendar & call for papers. IEEE Computational Science & Engineering, 2(4):93, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Eigenmann:1998:GEI

[EKP98] Rudolph Eigenmann, Laxmikant V. Kale, and David Padua. Guest Editor’s introduction: Languages for computational science and engineering. IEEE Computational Science & Engineering, 5(2):16–17, April/ June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2016.pdf.

Elliott:1995:CEEa

[Ell95a]  John Elliott.  CSE in Eu-

rope: Europort2: Increasing industrial confidence in parallel HPC. IEEE Computational Science & Engineering, 2(2):70–71, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Elliott:1995:CEEb

[Ell95b] John A. Elliott. CSE in Europe: Europort’s parallelized codes yield results. IEEE Computational Science & Engineering, 2 (4):19–21, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Ewing:1994:EPS

[EPWW94] Richard E. Ewing, Michael S. Pilant, J. Gordon Wade, and A. Ted Watson. Estimating parameters in scientific computation: a survey of experience from oil and groundwater modeling. IEEE Computational Science & Engineering, 1(3):19–31, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Ewald:1996:LCR

[Ewa96] Bob Ewald. Letters: Cray Research stands by its claim. IEEE Computational Science & Engineering, 3(4):95, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). The President of Cray Research comments on the dumping charges leveled against NEC in a proposed NCAR supercomputer acquisition. See [Smi96c] for an overview.

Fallah-Adl:1996:FAR

[FAJL^{+}96] Hassan Fallah-Adl, Joseph JáJá, Shunlin Liang, John Townshend, and Yoram J. Kaufman. Fast algorithms for removing atmospheric effects from satellite images. IEEE Computational Science & Engineering, 3(2): 66–77, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c2066abs.htm.

Fastook:1994:MIA

[Fas94] James L. Fastook. Modeling the ice age: The finite-element method in glaciology. IEEE Computational Science & Engineering, 1(1):55–67, Spring 1994. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Fortune:1995:WDI

[FGK^{+}95] Steven J. Fortune, David M. Gay, Brian W. Kernighan, Orlando Landron, Reinaldo A. Valenzuela, and Margaret H. Wright. WISE design of indoor wireless systems: Practical computation and optimization. IEEE Computational Science & Engineering, 2 (1):58–68, Spring 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c1058abs.htm.

Fine:1998:IMD

[FHST98] Steven S. Fine, Dongming Hwang, W. Ted Smith, and Timothy L. Turner. Improving model development with configuration management. IEEE Computational Science & Engineering, 5(1):56–65, January/ March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c1056.pdf;   http:
                                        //www.computer.org/cse/cs1998/
                                        c1056abs.htm.

Ferris:1995:CWF

[FJSP95] Michael Ferris and Jong-Shi-Pang. Conferences & workshops: First International Conference on Complementarity Problems; 11th ACM Symposium on Computational Geometry; High-Performance Computational Chemistry Tutorial and Workshop; 1995 IEEE Antennas and Propagation Society International Symposium; 1995 SIAM Annual Meeting. IEEE Computational Science & Engineering, 2 (4):90–92, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Feo:1998:CAT

[FKW98] John Feo, Simon Kahan, and Zhijun Wu. Crash analysis on the Tera MTA. IEEE Computational Science & Engineering, 5(4):53–59, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c4053.pdf;   http:
//www.computer.org/cse/cs1998/
c4053abs.htm.

Ford:1994:MSS

[FRN94] Ray Ford, Steven Running, and Ramakrishna Nemani. A modular system for scalable ecological modeling. IEEE Computational Science & Engineering, 1 (3):32–44, Fall 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Gallopoulos:1994:BN

[Gal94] Efstratios Gallopoulos. Book news. IEEE Computational Science & Engineering, 1(2):94, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Gallopoulos:1997:BNRb

[Gal97a] Efstratios Gallopoulos. Book news & reviews: Bookshelf.

IEEE Computational Science & Engineering, 4(3):83–84, July/ September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1997/pdf/c3083.
pdf.

Gallopoulos:1997:BNRa

[Gal97b] Stratis Gallopoulos. Book news & reviews: Bookshelf. IEEE Computational Science & Engineering, 4(1):83–84, January/ March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1083.pdf.

Gardner:1997:CEP

[Gar97] Henry F. Gardner. CSE in education: The puppy paradigm in computational science and engineering. IEEE Computational Science & Engineering, 4(3):4–6, 8, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c3004.pdf.

Gannon:1998:DCA

[GBS^{+}98] Dennis Gannon, Randall Bramley, Thomas Stuckey, Juan Villacis, Jayashree Balasubramanjian, Esra Akman, Fabian Breg, Shridhar Diwan, and Madhusudhan Govindaraju. Developing component architectures for distributed scientific problem solving. IEEE Computational Science & Engineering, 5(2):50–63, April/June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2050.pdf;   http:
//www.computer.org/cse/cs1998/
c2050abs.htm.

Germain:1997:FMU

[GCC97] Robert S. Germain, Andrea Califano, and Scott Colville. Fingerprint matching using transformation parameter clustering. IEEE Computational Science & Engineering, 4(4):42–49, October/December 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1997/pdf/c4042.
pdf;   http://www.computer.
org/cse/cs1998/c4042abs.htm.

Gao:1995:SSE

[GF95] Jiali Gao and Thomas R.

Furlani. Simulating solvent effects in organic chemistry: Combining quantum and molecular mechanics. IEEE Computational Science & Engineering, 2(3):24–33, Fall 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3024abs.htm.

Gallopoulos:1994:CTD

[GHR94] Efstratios Gallopoulos, Elias Houstis, and John R. Rice. Computer as thinker/doer: Problemsolving environments for computational science. IEEE Computational Science & Engineering, 1(2):11–23, Summer 1994. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Greengard:1998:AFM

[GHRW98] Leslie Greengard, Jingfang Huang, Vladimir Rokhlin, and Stephen Wandzura. Accelerating fast multipole methods for the Helmholtz equation at low frequencies. IEEE Computational Science & Engineering, 5(3):32–38, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c3032.pdf;   http:
                                        //www.computer.org/cse/cs1998/
                                        c3032abs.htm.

Gronbech-Jensen:1995:ACM

[GJGLB95] Niels Grønbech-Jensen, Tim Germann, Peter S. Lomdahl, and David M. Beazley. About the cover: Molecular reconnaissance. IEEE Computational Science & Engineering, 2(2):4–5, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Goudreau:1994:CSM

[Gou94] Gerald L. Goudreau. Computational structural mechanics: From national defense to national resource. IEEE Computational Science & Engineering, 1(1):33– 42, Spring 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Goodnick:1994:CWT

[GP94] Stephen M. Goodnick and Wolfgang Porod. Conferences & workshops: Third IWCE. IEEE Computational Science & Engineering, 1(3):92–93, Fall 1994. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Gueziec:1997:MIR

[GPA97] André P. Guéziec, Xavier Pennec, and Nicholas Ayache. Medical image registration using geometric hashing. IEEE Computational Science & Engineering, 4(4):29–41, October/December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c4029.pdf;   http:
//www.computer.org/cse/cs1998/
c4029abs.htm.

Graps:1995:IW

[Gra95] Amara Graps. An introduction to wavelets. IEEE Computational Science & Engineering, 2 (2):50–61, Summer 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c2050abs.htm.

Grosse:1996:NPC

[Gro96] Eric Grosse. Network programming and CSE. IEEE Computational Science & Engineering, 3(2):40–41, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). Discusses the emergence of systems like Java and Inferno for scientific visualization.

Gallopoulos:1997:OVC

[GS97] Efstratios Gallopoulos and Ahmed Sameh. One viewpoint on computational science: CSE: Content and product. IEEE Computational Science & Engineering, 4(2):39–43, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2039.pdf;   http:
//www.computer.org/cse/cs1998/
c2039abs.htm.

Gyure:1998:PMH

[GS98] Mark F. Gyure and Mark A. Stalzer. A prescription for the multilevel Helmholtz fast multipole method. IEEE Computational Science & Engineering, 5(3):39–47, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c3039.pdf;   http:
//www.computer.org/cse/cs1998/
c3039abs.htm.

Gueziec:1995:SRD

[Gué95] André Guéziec. Surface representation with deformable splines: Using decoupled variables. IEEE Computational Science & Engineering, 2(1): 69–80, Spring 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c1069abs.htm.

Gurd:1997:BNR

[Gur97] John R. Gurd. Book news & reviews: High Performance Computing: Challenges for Future Systems by David J. Kuck. IEEE Computational Science & Engineering, 4(1):85–86, January/ March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1083.pdf.

Gustafson:1997:CWD

[Gus97]  John Gustafson.  Confer-

ences & workshops: Distributed means clusters, networks — and NT? IEEE Computational Science & Engineering, 4(3):87, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c3086.pdf. Review

of the Sixth High Performance Distributed Computing Conference held at Portland State University in August 1997.

Gustafson:1998:CVF

[Gus98]  John Gustafson.  Computa-

tional verifiability and feasibility of the ASCI program. IEEE Computational Science & Engineering, 5(1):36–45, January/ March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c1036.pdf;   http:
//www.computer.org/cse/cs1998/

c1036abs.htm. Discusses recent progress in interval arithmetic and its relevance to error estimation in very large computations of the type envisioned for the US ASCI (Advanced Strategic Computing Initiative) project.

Gothard:1994:CMT

[GVR94] Griffin K. Gothard, Douglas A. Vechinski, and Sadasiva M. Rao. Computational methods in transient electromagnetics: a selective survey. IEEE Computational Science & Engineering, 1(2):50– 59, Summer 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Greengard:1998:GEI

[GW98] Leslie Greengard and Stephen Wandzura. Guest Editor’s introduction: Fast multipole methods. IEEE Computational Science & Engineering, 5(3):16–18, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c3016.pdf.

Haggerty:1994:ACC

[Hag94] Michael Haggerty. About the cover: Catching Einstein’s waves. IEEE Computational Science & Engineering, 1(1):8–9, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Hanson:1995:BNR

[Han95] Andrew J. Hanson. Book news & reviews: Scientific Visualization: Advances and Challenges, edited by Lawrence J. Rosenblum et al. IEEE Computational Science & Engineering, 2(4):87– 88, Winter 1995. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Hatton:1997:EES

[Hat97] Les Hatton. The T experiments:

Errors in scientific software. IEEE Computational Science & Engineering, 4(2):27–38, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2027.pdf;   http:
//www.computer.org/cse/cs1998/
c2027abs.htm.

Heath:1998:VRR

[HD98] Michael T. Heath and William A.

Dick. Virtual rocketry: Rocket science meets computer science. IEEE Computational Science & Engineering, 5(1):16– 26, January/March 1998. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1998/pdf/c1016.
pdf;   http://www.computer.
org/cse/cs1998/c1016abs.htm.

Hwang:1995:PMD

[HDS^{+}95] Yuan-Shin Hwang, Raja Das, Joel H. Saltz, Milan Hodoscek, and Bernard R. Brooks. Parallelizing molecular dynamics programs for distributed-memory machines. IEEE Computational Science & Engineering, 2(2):

18–29, Summer 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c2018abs.htm.

Houstis:1997:GEI

[HGBR97] Elias Houstis, Efstratios Gallopoulos, Randall Bramley, and John Rice. Guest Editors’ introduction: Problem-solving environments for computational science. IEEE Computational Science & Engineering, 4(3):18–21, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c3018.pdf;   http:
                                        //www.computer.org/cse/cs1998/
                                        c3018abs.htm.

Hoekstra:1996:ACO

[HHS96] A. Hoekstra, B. Hertzberger, and P. Sloot. About the cover: The optical “fingerprints” of cells: Catching the bad guys. IEEE Computational Science & Engineering, 3(1):4–5, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Higham:1998:BNR

[Hig98] Nicholas J. Higham. Book news & reviews: Review of “Christoph W. Ueberhuber, Numerical Computation: Methods, Software and Analysis, Springer-Verlag, 1997”. IEEE Computational Science & Engineering, 5(1):79, January/ March 1998. CODEN ISCEE4.

ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c1079.pdf.

Houstis:1996:BNR

[Hou96] Elias N. Houstis. Book news & reviews: Designing and Building Parallel Programs by Ian T. Foster. IEEE Computational Science & Engineering, 3(4):87, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Hussaini:1995:SRI

[Hus95] M. Y. Hussaini. Site report:

ICASE: NASA Langley’s CSE center. IEEE Computational Science & Engineering, 2(1):6–14, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Issaevitch:1996:WSC

[Iss96] Tom Issaevitch. What should computer scientists teach to physical scientists and engineers? part 2. response to Wilson: forget multiple tools; use Mathematica. IEEE Computational Science & Engineering, 3(3):59–61, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3059abs.htm.

Johnson:1994:CES

[JA94] Christopher R. Johnson and Peter Alfeld. Computational engineering and science at the University of Utah. IEEE Computational Science & Engineering, 1 (3):7–9, Fall 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Jalby:1994:CEE

[Jal94] William Jalby. CSE in Europe: Europort: Building confidence in parallel HPC. IEEE Computational Science & Engineering, 1 (4):75, Winter 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Jessup:1996:WSC

[JG96] Elizabeth R. Jessup and Roscoe C. Giles. What should computer scientists teach to physical scientists and engineers? part 2. response to Wilson: Teach computing in context. IEEE Computational Science & Engineering, 3(3):54–58, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3054abs.htm.

Proposes that Java be included in a week-long course of “computer science for physical scientists.”.

Johnston:1995:PPJ

[JiI95] Scott C. Johnston and Ken ichi Imai. Parallel processing in Japan: National and corporate trends. IEEE Computational Science & Engineering, 2 (1):32–41, Spring 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c1032abs.htm.

Johnson:1995:TEI

[Joh95] Christopher R. Johnson. Theme Editor’s introduction: Computational inverse problems in medicine. IEEE Computational Science & Engineering, 2 (4):42–45, Winter 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c4042abs.htm.

Johnson:1996:CWS

[Joh96] Christopher R. Johnson. Conferences & workshops: SciTools ’96. IEEE Computational Science & Engineering, 3(4):83–84, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Jones:1996:TSL

[Jon96] Anita K. Jones. Taking stock, looking ahead: Part 2. modernizing high-performance computing for the military. IEEE Computational Science & Engineering, 3(3):71–74, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3071abs.htm.

Jalby:1994:CEA

[JW94] William Jalby and Harry A. G.

Wijshoff. CSE in Europe: The APPARC Project. IEEE Computational Science & Engineering, 1 (2):83, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Jalby:1995:CE

[JW95] William Jalby and Harry A. G. Wijshoff. CSE in Europe: More on Europort-1. IEEE Computational Science & Engineering, 2 (3):86–87, Fall 1995. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Joshi:1996:NFS

[JWR^{+}96] Anupam Joshi, Sanjiva Weerawarana, Narendran Ramakrishnan, Elias N. Houstis, and John R. Rice. Neuro-fuzzy support for problem-solving environments: a step toward automated solution of PDEs. IEEE Computational Science & Engineering, 3(1):44–56, Spring 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c1044abs.htm.

Jordan:1996:PIV

[JYR^{+}96] Kirk E. Jordan, David A. Yuen, David M. Reuteler, Shuxia Zhang, and Robert Haimes. Parallel interactive visualization of 3D mantle convection. IEEE Computational Science & Engineering, 3(4): 29–37, Winter 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c4029abs.htm.

Kahaner:1994:CAE

[Kah94a] David K. Kahaner. CSE in Asia: Emphasizing applications: Inverse problem-solving; HPC in Asia; simulation in industry; automating vehicle navigation. IEEE Computational Science & Engineering, 1(4):73–74, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Kahaner:1994:CJS

[Kah94b] David K. Kahaner. CSE in Japan: Stiff competition: Highperformance computing in Japan. IEEE Computational Science & Engineering, 1(2):84–86, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Kahaner:1995:CAW

[Kah95] David K. Kahaner. CSE in Asia:

a warming R&D climate. IEEE Computational Science & Engineering, 2(2):94, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Kalos:1995:EPG

[Kal95] Malvin H. Kalos. Editor’s perspective: Globally scalable computing. IEEE Computational Science & Engineering, 2(2):96, 71, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Kale:1998:PLC

[Kal98]  Laxmikant V. Kale.  Pro-

gramming languages for CSE: The state of the art. IEEE Computational Science & Engineering, 5(2):18–26, April/ June 1998. CODEN ISCEE4.

ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2018.pdf;   http:
//www.computer.org/cse/cs1998/
c2018abs.htm.

Kincaid:1998:CWI

[KE98] David R. Kincaid and Anne C. Elster. Conferences & workshops: Iterative Methods Symposium Honors David M. Young, Jr. IEEE Computational Science & Engineering, 5(4):12–15, October/December 1998. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1998/pdf/c4011.
pdf.

King:1995:NJE

[Kin95]  Scott D. King.  A numeri-

cal journey to the earth’s interior. IEEE Computational Science & Engineering, 2(3): 12–23, Fall 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3012abs.htm.

Knox:1997:PSW

[KKLK97] Robert G. Knox, Virginia L. Kalb, Elissa R. Levine, and David J. Kendig. A problemsolving workbench for interactive simulation of ecosystems. IEEE Computational Science & Engineering, 4(3):52– 60, July/September 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1997/pdf/c3052.
pdf;   http://www.computer.
org/cse/cs1998/c3052abs.htm.

Kapur:1998:IEE

[KL98] Sharad Kapur and David E.

Long. IES 3: Efficient electrostatic and electromagnetic simulation. IEEE Computational Science & Engineering, 5(4):60–67, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c4060.pdf;   http:
//www.computer.org/cse/cs1998/
c4060abs.htm.

Koebel:1998:CWC

[Koe98] Chuck Koebel. Conferences & workshops: Conferences for Scientific Applications. IEEE Computational Science & Engineering, 5(3):91–94, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c3091.pdf.

Kuehlmann:1998:CWI

[Kue98]  Andreas Kuehlmann.  Con-

ferences & workshops: International Conference on Computer Design ’98. IEEE Computational Science & Engineering, 5(4):11, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c4011.pdf.

Landau:1995:BNR

[Lan95] Rubin H. Landau. Book news & reviews: Computation in Modern Physics by William R. Gibbs. IEEE Computational Science & Engineering, 2(3):84–85, Fall 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Landau:1996:RCS

[Lan96] Rubin H. Landau. Response: Computer scientists should not teach computational science. IEEE Computational Science & Engineering, 3(2):55–61, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://www.

                                        computer.org/cse/cs1998/c2055abs.
                                        htm.

Larzelere:1998:CSC

[Lar98] Alex R. Larzelere, II. Creating simulation capabilities. IEEE Computational Science & Engineering, 5(1):27–35, January/ March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c1027.pdf;   http:
                                        //www.computer.org/cse/cs1998/
                                        c1027abs.htm.

Lehoucq:1995:CWB

[Leh95] Richard B. Lehoucq. Conferences & workshops: 16th Biennial on Numerical Analysis; Principles and Practices of Parallel Programming & User Interface Workshop on Parallel; Programming Software; ICIAM; Computing for the Social Sciences; Numerical Simulation in the Environmental and Earth Sciences. IEEE Computational Science & Engineering, 2(3):74–76, Fall 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Levine:1997:SIS

[LFH^{+}97] David Levine, Michael Facello, Phillip Hallstrom, Gregory Reeder, Brian Walenz, and Fred Stevens. Stalk: An interactive system for virtual molecular docking. IEEE Computational Science & Engineering, 4(2):55–65, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2055.pdf;   http:
//www.computer.org/cse/cs1998/
c2055abs.htm.

Lee:1994:IPP

[LG94] Peter ChiKeung Lee and Sumit Ghosh. International payments processing in real time: a distributed architecture. IEEE Computational Science & Engineering, 1(3):65–79, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Lee:1996:SAD

[LG96] Tony Lee and Sumit Ghosh.

Simulating asynchronous, decentralized military command and control. IEEE Computational Science & Engineering, 3 (4):69–79, Winter 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c4069abs.htm.

Lanczycki:1998:PCS

[LJT^{+}98] Christopher J. Lanczycki, Calvin A. Johnson, Benes L. Trus, James F. Conway, Alasdair C. Steven, and Robert L. Martino. Parallel computing strategies for determining viral capsid structure by cryo-electron microscopy. IEEE Computational Science & Engineering, 5(2):76–91, April/ June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c2076.pdf;   http:
                                        //www.computer.org/cse/cs1998/
                                        c2076abs.htm.

Leech:1996:SVS

[LPH96] Jonathan Leech, Jan F. Prins, and Jan Hermans. SMD: Visual steering of molecular dynamics for protein design. IEEE Computational Science & Engineering, 3(4):38–45, Winter 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c4038abs.htm.

Marchioro:1996:CEC

[Mar96] Thomas L. Marchioro, II. CSE in education: CSE educators win DoE awards. IEEE Computational Science & Engineering, 3 (4):91–92, 94, Winter 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Markus:1997:DIM

[Mar97]  Karen W. Markus.  Devel-

oping infrastructure to massproduce MEMS. IEEE Computational Science & Engineering, 4(1):49–54, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1049.pdf;   http:
//www.computer.org/cse/cs1998/
c1049abs.htm.

Marchioro:1998:CEL

[Mar98a] Tom Marchioro. CSE education: Larry Geller: a CSE prescription for the pharmaceutical industry: 1997 computational science education awards. IEEE Computational Science & Engineering, 5(2):10–15, April/ June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2010.pdf.

Marchioro:1998:NAW

[Mar98b] Tom Marchioro. News analysis: The wind and the lion. IEEE Computational Science & Engineering, 5(3):4–7, July/September 1998. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1998/pdf/c3004.
pdf.

Margenov:1998:BNR

[Mar98c] Svetozar Margenov. Book news & reviews: Introduction to Fortran 90 for Engineers and Scientists. IEEE Computational Science & Engineering, 5(3):87, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c3087.pdf.

Mascagni:1998:CWH

[Mas98] Michael Mascagni. Conferences & workshops: High-Performance Monte Carlo Tools. IEEE Computational Science & Engineering, 5(2):97–98, April/ June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2097.pdf.

McConnell:1996:RTP

[McC96] Steve McConnell. Response: Teach programming principles, not “tools and tips”. IEEE Computational Science & Engineering, 3(2):62–65, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://www.

computer.org/cse/cs1998/c2062abs.
htm.

McMullen:1997:TNR

[McM97] Donald F. McMullen. Technology news & reviews: Sun’s promising move toward HPC. IEEE Computational Science & Engineering, 4(2):78–81, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2078.pdf.

McQueeney:1996:WSC

[McQ96] David F. McQueeney. What should computer scientists teach to physical scientists and engineers? part 2. response: a week won’t do it. IEEE Computational Science & Engineering, 3(3):58–59, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3058abs.htm.

Meyers:1994:AAK

[Mey94] Ware Meyers. About the articles: Kenneth G. Wilson: From quantum chromodynamics to redesigning education. IEEE Computational Science & Engineering, 1(4):14, Winter 1994. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Miller:1994:BNR

[Mil94] Edmund K. Miller. Book news & reviews: Numerical Recipes: The Art of Scientific Computing, 2nd Edition, by William H. Press, Saul A. Teukolsky, William T. Vetterling, and Brian P. Flannery. IEEE Computational Science & Engineering, 1(3):85–86, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Minty:1998:CEF

[Min98] Elspeth M. Minty. CSE in education: First 10 years of the EPCC summer scholarship programme.

IEEE Computational Science & Engineering, 5(1):6–9, January/ March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c1006.pdf.

Munthe-Kaas:1996:CWS

[MK96] Hans Munthe-Kaas. Conferences & workshops: State of the Art in Numerical Analysis. IEEE Computational Science & Engineering, 3(2):90–91, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Marchioro:1997:WBE

[ML97] Thomas L. Marchioro, II and Rubin H. Landau. Web-based education in computational science and engineering. IEEE Computational Science & Engineering, 4(2):19–26, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2019.pdf;   http:
//www.computer.org/cse/cs1998/
c2019abs.htm.

Martin:1998:CCD

[MM98a] Ioana M. Martin and Dan C. Marinescu. Concurrent computation and data visualization for spherical-virus structure determination. IEEE Computational Science & Engineering, 5(4):40–52, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c4040.pdf;   http:
//www.computer.org/cse/cs1998/
c4040abs.htm.

Moreira:1998:FCC

[MM98b] José E. Moreira and Samuel P.

Midkiff. Fortran 90 in CSE: a case study. IEEE Computational Science & Engineering, 5(2):39–49, April/June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2039.pdf;   http:
//www.computer.org/cse/cs1998/
c2039abs.htm.

Marchioro:1995:CEU

[MMP95] Thomas L. Marchioro, II, David M.

Martin, and W. Donald Payne. CSE education: UCES: An undergraduate CSE initiative. IEEE Computational Science & Engineering, 2(3):69–73, Fall 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (elec-

tronic).  URL http://www.
computer.org/cse/cs1998/c3069abs.
htm.

Mohr:1996:NAU

[Moh96] Jeffrey Mohr. News analysis:

Upsizing, downsizing, rightsizing, wrongsizing. IEEE Computational Science & Engineering, 3 (2):99–103, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Mokkarala:1997:LE

[Mok97] Viggy Mokkarala. Letter to the Editor. IEEE Computational Science & Engineering, 4(1):6, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c1001.pdf;   http:
                                        //www.sgi.com/techpubs/lib/
                                        makepage.cgi?007-3430-001.

Comments on the measured speedups of various parallel programs on SGI Origin 2000 systems.

Moreira:1996:BNR

[Mor96] José E. Moreira. Book news & reviews: High Performance Computing Demystified by David Loshin. IEEE Computational Science & Engineering, 3(1):86– 87, Spring 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Morjaria:1996:CWM

[MS96] Mahesh Morjaria and Fadil Santosa. CSE at work: Monitoring complex systems with causal networks. IEEE Computational Science & Engineering, 3(4):9– 10, Winter 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Miller:1994:PCB

[MSH94] Webb Miller, Scott Schwartz, and Ross C. Hardison. A point of contact between computer science and molecular biology. IEEE Computational Science & Engineering, 1(1):69–78, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Muraoka:1996:HCA

[Mur96] Yoichi Muraoka. How to cooperate across the ocean. IEEE Computational Science & Engineering, 3(2):45, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Marefat:1997:IVA

[MVY97] Michael M. Marefat, Andrew F.

Varecka, and Jeff Yost. An intelligent visualization agent for simulation-based decision support. IEEE Computational Science & Engineering, 4(3):72–82, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c3072.pdf;   http:
//www.computer.org/cse/cs1998/
c3072abs.htm.

Mehrotra:1998:LSM

[MVZ98] Piyush Mehrotra, John Van Rosendale, and Hans P. Zima. Language support for multidisciplinary applications. IEEE Computational Science & Engineering, 5(2):64–75, April/ June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2064.pdf;   http:
//www.computer.org/cse/cs1998/
c2064abs.htm.

Markus:1997:SCW

[MWHR97] Shahani Markus, Sanjiva Weerawarana, Elias N. Houstis, and John R. Rice. Scientific computing via the Web: The Net Pellpack PSE server. IEEE Computational Science & Engineering, 4(3):43–51, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c3043.pdf;   http:
                                        //www.computer.org/cse/cs1998/
                                        c3043abs.htm.

Nakano:1998:ASN

[NBC^{+}98] Aiichiro Nakano, Martina E. Bachlechner, Timothy J. Campbell, Rajiv K. Kalia, Andrey Omeltchenko, Kenji Tsuruta, and Priya Vashista. Atomistic simulation of nanostructured materials. IEEE Computational Science & Engineering, 5(4):68–78, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c4068.pdf;   http:
                                        //www.computer.org/cse/cs1998/
                                        c4068abs.htm.

Noor:1995:SRU

[Noo95] Ahmed K. Noor. Site report: UVA Center for Computational Structures Technology. IEEE Computational Science & Engineering, 2(4):10–14, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Oden:1995:EPC

[Ode95] J. Tinsley Oden. Educational programs in CSE; UT Austin establishes CAM program. IEEE Computational Science & Engineering, 2(2):7–9, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Obeysekare:1997:VID

[OGP97] Upul Obeysekare, Fernando F.

Grinstein, and Gopal Patnaik. The visual interactive desktop laboratory. IEEE Computational Science & Engineering, 4(1):63–71, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1063.pdf;   http:
//www.computer.org/cse/cs1998/
c1063abs.htm.

OLeary:1994:EPP

[O’L94] Dianne P. O’Leary. From an Editor’s perspective: Parallel computing emerging from a time warp. IEEE Computational Science & Engineering, 1(4):1, 15, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

OLeary:1997:TCS

[O’L97]  Dianne P. O’Leary.  Team-

work: Computational science and applied mathematics. IEEE Computational Science & Engineering, 4(2):13–18, April/ June 1997. CODEN ISCEE4.

ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c2013.pdf;   http:
                                        //www.computer.org/cse/cs1998/
                                        c2013abs.htm.

Padua:1996:TSL

[Pad96] David A. Padua. Taking stock, looking ahead: Part 2. outline of a roadmap for compiler technology. IEEE Computational Science & Engineering, 3(3):65–66, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c3065abs.htm.

Panoff:1995:CEF

[Pan95] Robert M. Panoff. CSE education: The four A’s of CSE education. IEEE Computational Science & Engineering, 2(4):6– 7, 9, Winter 1995. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Pancake:1996:PY

[Pan96] Cherri M. Pancake. Is parallelism for you? IEEE Computational Science & Engineering, 3(2):18–37, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c2018abs.htm.

Papakhian:1998:TNR

[Pap98]  Mary Papakhian.  Technol-

ogy news & reviews: Comparing job-management systems: the user’s perspective. IEEE Computational Science & Engineering, 5(2):4–9, April/June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2004.pdf.

Patt:1996:MSC

[Pat96] Yale N. Patt. The microprocessor for scientific computing in the year 2000. IEEE Computational Science & Engineering, 3(2):42– 43, Summer 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Pryor:1998:CWS

[PC98] Daniel Pryor and Steve Cuccaro. Conferences & workshops: SC ’97: Smart New World; ICIP ’97 — Borrowing from Other Fields. IEEE Computational Science & Engineering, 5(1):80–83, January/ March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c1080.pdf.

Peeters:1997:CCM

[Pee97]  Eric Peeters.  Challenges in

commercializing MEMS. IEEE Computational Science & Engineering, 4(1):44–48, January/ March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1044.pdf;   http:
//www.computer.org/cse/cs1998/
c1044abs.htm.

Pehle:1995:CWC

[Peh95] Hans Joachim Pehle. CSE at work: Crunching numbers, shaping steel. IEEE Computational Science & Engineering, 2(4):15– 18, Winter 1995. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Peters:1995:CDL

[PFT^{+}95] Gunther H. Peters, Thomas Frimurer, Søren Toxvaerd, Ole H. Olsen, and Allan Svendsen. Computing at different levels of approximation: Examples in molecular biology. IEEE Computational Science & Engineering, 2(3):43–54, Fall 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c3043abs.htm.

Philippe:1996:BNR

[Phi96] Bernard Philippe. Book news & reviews: Introduction to the Numerical Solution of Markov Chains by W. J. Stewart. IEEE Computational Science & Engineering, 3(2):94–95, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Parker:1997:CSS

[PJB97] Steven G. Parker, Christopher R. Johnson, and David Beazley. Computational steering software systems and strategies. IEEE Computational Science & Engineering, 4(4):50–59, October/December 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1997/pdf/c4050.
pdf;   http://www.computer.
org/cse/cs1998/c4050abs.htm.

Plemmons:1997:BNR

[Ple97] Robert J. Plemmons. Book news & reviews: Imaging Through Turbulence by M. Roggemann and B. Welsh. IEEE Computational Science & Engineering, 4(3):85, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c3083.pdf.

Payne:1995:CEA

[PM95] W. Donald Payne and Thomas L.

Marchioro, II. CSE education: Awards recognize excellence and innovation in CSE education. IEEE Computational Science & Engineering, 2(4):8– 9, Winter 1995. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Plummer:1998:CEC

[PM98] Martin Plummer and Thomas L.

Marchioro, II. CSE education: Canadian efforts grow. IEEE Computational Science & Engineering, 5(3):11–15, July/ September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1998/pdf/c3011.
pdf.

Prince:1995:LEF

[PRSE95] Tim Prince, Dietrich Rhein, Ahmed Sameh, and Alexander Elder. Letters to the Editor: Feedback. IEEE Computational Science & Engineering, 2(2):6, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Pryor:1997:CWF

[Pry97] Dan Pryor. Conferences & workshops: Frontiers ’96. IEEE Computational Science & Engineering, 4(1):101–102, January/ March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c1098.pdf.

Pierce:1995:ACV

[PS95] David Pierce and Benjamin K. Stephens. About the cover: Visualizing geophysical data: Teasing meaning from models. IEEE Computational Science & Engineering, 2(4):4–5, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Quarteroni:1996:SRC

[QAF^{+}96] Alfio Quarteroni, Davide Ambrosi, Luca Formaggia, Giorgio Fotia, Marco Manzine, Marco Mulas, and Luca Stolcis. Site report: Computational fluid dynamics at CRS4, Italy. IEEE Computational Science & Engineering, 3(3):4–8, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Rice:1996:SSL

[RB96] John R. Rice and Ronald F.

Boisvert. From scientific software libraries to problem-solving environments. IEEE Computational Science & Engineering, 3(3):44–53, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3044abs.htm.

Rigoutsos:1994:SPS

[RC94] Isidore Rigoutsos and Andrea Califano. Searching in parallel for similar strings. IEEE Computational Science & Engineering, 1 (2):60–75, Summer 1994. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Rice:1994:APC

[Ric94a] John R. Rice. Academic programs in computational science and engineering. IEEE Computational Science & Engineering, 1 (1):13–21, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Rice:1994:EPW

[Ric94b] John R. Rice. Editor’s perspectives: Whither computational science and engineering. IEEE Computational Science & Engineering, 1(2):96, Summer 1994. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Rice:1995:CSF

[Ric95] John R. Rice. Computational science and the future of computing research (linked to full text). IEEE Computational Science & Engineering, 2(4): 35–41, Winter 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c4035abs.htm.

Rice:1996:CWA

[Ric96a] John R. Rice. Conferences & workshops: Announcing the first IEEE-CS Workshop on Computational Science and Engineering, Purdue University, Indiana, October 21–26, 1996. IEEE Computational Science & Engineering, 3 (1):88, Spring 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Rice:1996:CWF

[Ric96b] John R. Rice. Conferences & workshops: First IEEE Computer Society Workshop on CSE. IEEE Computational Science & Engineering, 3(4):86, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Rice:1997:CCC

[Ric97a] John R. Rice. The CSE community converges: First IEEE computer society workshop on computational science and engineering. IEEE Computational Science & Engineering, 4(2):10–11, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2010.pdf;   http:

//www.computer.org/cse/cs1998/
c2010abs.htm.

Rice:1997:FSS

[Ric97b] John R. Rice. Future scientific software systems. IEEE Computational Science & Engineering, 4(2):44–48, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2044.pdf;   http:
//www.computer.org/cse/cs1998/
c2044abs.htm.

Riganati:1994:EPR

[Rig94a] John P. Riganati. Editors’ perspectives: Reality check: Systems and performance. IEEE Computational Science & Engineering, 1 (3):96, Fall 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Riganati:1994:EPC

[Rig94b] John P. Riganati. Educational programs in CSE: Are you interested in a technical committee on CSE education? IEEE Computational Science & Engineering, 1 (2):76, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Rigoutsos:1997:GEI

[RW97] Isidore Rigoutsos and Haim J.

Wolfson. Guest Editors’ introduction: Geometric hashing. IEEE Computational Science & Engineering, 4(4):9, October/December 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://dlib.computer.
                                        org/cs/books/cs1997/pdf/c4009.
                                        pdf.

Saad:1995:BNR

[Saa95] Youcef Saad. Book news & reviews: Concurrent Scientific Computing by Eric F. Van de Velde. IEEE Computational Science & Engineering, 2(1):94, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Saied:1996:BNR

[Sai96] Faisal Saied. Book news & reviews: An Introduction to Computational Science and Mathematics by Charles Van Loan. IEEE Computational Science & Engineering, 3(4):89, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Sameh:1994:AA

[Sam94a] Ahmed Sameh. About the articles. IEEE Computational Science & Engineering, 1(2):9, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Sameh:1994:AAW

[Sam94b] Ahmed Sameh. About the articles: What if? IEEE Computational Science & Engineering, 1 (3):15–16, Fall 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Sameh:1994:EC

[Sam94c] Ahmed Sameh. From the Editorin-Chief. IEEE Computational Science & Engineering, 1(1):1, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Sameh:1995:ECA

[Sam95a] Ahmed Sameh. From the Editorin-Chief: An auspicious beginning. IEEE Computational Science & Engineering, 2(1):1, 96, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Sameh:1995:ECF

[Sam95b] Ahmed Sameh. From the Editorin-Chief: Farewell, CS&E, fare well, CSE. IEEE Computational Science & Engineering, 2(4):1, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Sameh:1995:ECL

[Sam95c] Ahmed Sameh. From the Editorin-Chief: Looking back, looking forward: Evaluation and vision. IEEE Computational Science & Engineering, 2(2):1, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Santosa:1996:CWR

[San96] Fadil Santosa. CSE at work:

Real CSE problems from industry. IEEE Computational Science & Engineering, 3(3):76–77, Fall 1996. CODEN ISCEE4.

ISSN 1070-9924 (print), 1558-190X (electronic).

Senturia:1997:SBM

[SAW97] Stephen D. Senturia, Narayan Aluru, and Jacob White. Simulating the behavior of MEMS devices: Computational methods and needs. IEEE Computational Science & Engineering, 4(1):30–43, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1030.pdf;   http:
//www.computer.org/cse/cs1998/
c1030abs.htm.

Stiles:1994:ACW

[SB94] Joel Stiles and Thomas Bartol, Jr. About the cover: Walking randomly in a synapse. IEEE Computational Science & Engineering, 1(3):4–5, Fall 1994. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Schlick:1996:CWM

[SB96] Tamar Schlick and Achi Brandt. Conferences & workshops: A Multigrid Tutorial with Applications to Molecular Dynamics. IEEE Computational Science & Engineering, 3(3):78–79, 82, Fall 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Song:1998:FIS

[SC98] Jiming Song and Weng Cho Chew. The Fast Illinois Solver Code: Requirements and scaling properties. IEEE Computational Science & Engineering, 5(3):19–23, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c3019.pdf;   http:
//www.computer.org/cse/cs1998/
c3019abs.htm.

Schneider:1995:BNR

[Sch95a] David J. Schneider. Book news & reviews: Parallel Computing Works! IEEE Computational Science & Engineering, 2(2):91– 92, Summer 1995. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Schulz:1995:CEP

[Sch95b] Uwe Schulz. CSE in Europe: The Permpar Project. IEEE Computational Science & Engineering, 2 (1):88–89, Spring 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Santosa:1996:CWW

[SGS^{+}96] Fadil Santosa, Myron Ginsberg, Danny Sorensen, Francis Sullivan, Laxmikant Kalé, Kirk Jordan, and Nick Camp. CSE at work: What industry needs from academia: CSE education in the 21st Century. IEEE Computational Science & Engineering, 3 (2):4–10, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Shankar:1998:QTC

[SHR98] Vijaya Shankar, William F. Hall, and S. V. Ramakrishnan. Quick-turnaround computational simulation. IEEE Computational Science & Engineering, 5(1):66–78, January/ March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c1066.pdf;   http:
                                        //www.computer.org/cse/cs1998/
                                        c1066abs.htm.

Simon:1997:SRR

[Sim97] Horst D. Simon. Site report: Reinventing the supercomputer center at NERSC. IEEE Computational Science & Engineering, 4(3):14–17, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c3014.pdf.

Silva:1996:PHP

[SKP96] Cláudio T. Silva, Arie E. Kaufman, and Constantine Pavlakos. PVR: High-performance volume rendering. IEEE Computational Science & Engineering, 3 (4):18–28, Winter 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c4018abs.htm.

Smedley:1998:CWI

[Sme98]  Trevor J. Smedley.  Con-

ferences & workshops: 14th IEEE Symposium on Visual Languages. IEEE Computational Science & Engineering, 5(4):11–12, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c4011.pdf.

Smith:1995:ICS

[Smi95a] Norris Parker Smith. Interfaces: Cray/SGI matchup adds drama as Kiwis win America’s Cup. IEEE Computational Science & Engineering, 2(2):72–73, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Smith:1995:IMN

[Smi95b] Norris Parker Smith. Interfaces: Midsummer night’s nightmare; teraflops rex; hierarchical archaeology. IEEE Computational Science & Engineering, 2 (3):88, 87, Fall 1995. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Smith:1996:IHH

[Smi96a] Norris Parker Smith. Interfaces:

Hardware for high-performance computing: Abstract progress, painful consolidation. IEEE Computational Science & Engineering, 3(2):104, 103, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Smith:1996:IHC

[Smi96b] Norris Parker Smith. Interfaces:

HPC and CSE in Japan: a new spirit? IEEE Computational Science & Engineering, 3(4):96, 95, Winter 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Smith:1996:IRH

[Smi96c] Norris Parker Smith. Interfaces: Reciprocity in high-performance trade. IEEE Computational Science & Engineering, 3(3):88, 87, Fall 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). See letter from Cray Research President [Ewa96].

Smith:1996:ITM

[Smi96d] Norris Parker Smith. Interfaces: Technoculture: Moving from rejection to accommodation. IEEE Computational Science & Engineering, 3(1):96, 91, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Smith:1997:IRH

[Smi97a] Norris Parker Smith. Interfaces: Rescuing humanity from the compugotcha. IEEE Computational Science & Engineering, 4(3):96, 95, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c3096.pdf.

Smith:1997:INO

[Smi97b] Norris Parker Smith. Interfaces: The new-order NSF centers: What will “partnership” and “infrastructure” really mean?

IEEE Computational Science & Engineering, 4(2):88, 87, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2088.pdf.

Smith:1997:IWS

[Smi97c] Norris Parker Smith. Interfaces:

What shapes computing’s future? more than hardware. IEEE Computational Science & Engineering, 4(1):104, 103, January/ March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1104.pdf.

Smith:1998:IEA

[Smi98a] Norris Parker Smith. Interfaces: Euphoritopia, the anxiety deficit, and econoflu. IEEE Computational Science & Engineering, 5(3):96–97, July/September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c3096.pdf.

Smith:1998:INL

[Smi98b] Norris Parker Smith. Interfaces:

Has El Niño let us down? IEEE Computational Science & Engineering, 5(1):88, 87, January/ March 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c1088.pdf.

Smith:1998:IMC

[Smi98c] Norris Parker Smith. Interfaces: The millennium is close at hand: Stock up on zeroes. IEEE Computational Science & Engineering, 5(2):104, 103, April/ June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c2104.pdf.

Smith:1998:ITT

[Smi98d] Norris Parker Smith. Interfaces: The turmoil turmoil and computational finance. IEEE Computational Science & Engineering, 5(4):88–90, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c4088.pdf.

Song:1994:LLE

[SN94] Deyang Song and Michael L. Norman. Looking in, looking out: Exploring multiscale data with virtual reality. IEEE Computational Science & Engineering, 1 (3):53–64, Fall 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Solomon:1996:CWN

[Sol96]  Stephen D. Solomon.  CSE

at work: a new market for CSE: An engineer goes to Wall Street. IEEE Computational Science & Engineering, 3(1):6– 7, Spring 1996. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Siegel:1996:TSL

[SS96] Howard Jay Siegel and Craig B.

Stunkel. Taking stock, looking ahead: Part 2. inside parallel computers: Trends in interconnection networks. IEEE Computational Science & Engineering, 3(3):69–71, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3069abs.htm.

Stevenson:1997:HGC

[Ste97] D. E. Stevenson. How goes CSE? thoughts on the IEEE CS workshop at purdue. IEEE Computational Science & Engineering, 4(2):49–54, April/ June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2049.pdf;   http:
//www.computer.org/cse/cs1998/
c2049abs.htm.

Spezzano:1996:PCT

[STG^{+}96] Giandomenico Spezzano, Domenico Talia, Salvatore Di Gregorio, Rocco Rongo, and William Spataro. A parallel cellular tool for interactive modeling and simulation. IEEE Computational Science & Engineering, 3(3):33–43, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3033abs.htm.

Stonebraker:1994:SRF

[Sto94] Michael Stonebraker. Sequoia 2000: a reflection on the first three years. IEEE Computational Science & Engineering, 1(4):63– 72, Winter 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Strok:1994:MEH

[Str94a] Dale C. Strok. From the Managing Editor: How are we doing? IEEE Computational Science & Engineering, 1(3):1, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Strok:1994:NJI

[Str94b] Dale C. Strok. In the news: Jupiter impacts: Resolution makes a big difference. supercomputer farming down under. HPF Forum welcomes comments. Smithsonian Awards honor computational scientists. low-life computer viruses. PVM developers get R&D-100 award. the eyes have it. neural nets detect breast cancer. better cars through cooperation. parallel version of global climate model. Lockheed to run Idaho National Engineering Lab. public-private partners: new drugs, new software. IEEE Computational Science & Engineering, 1(3):88–90, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Strok:1994:NWG

[Str94c] Dale C. Strok. In the news: What’s so good about chaos? making planes safer and more efficient. FMC honored for industrial research. there’s “gold” in them thar’ labs. … why do riblets work? Computer Society on-line. electromagnetics on line. finding millisecond pulsars. IEEE Computational Science & Engineering, 1(2):87–88, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Sullivan:1994:AEC

[Sul94] Francis Sullivan. From the Associate Editor-in-Chief: Cautionary note. IEEE Computational Science & Engineering, 1(2):1, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Sullivan:1995:AEC

[Sul95] Francis Sullivan. From the Associate Editor-in-Chief: Who needs efficient algorithms? IEEE Computational Science & Engineering, 2(3):1, Fall 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Sullivan:1997:AEC

[Sul97] Francis Sullivan. From the Associate Editor-in-Chief: Will the circle be unbroken? (A meditation on the passing of Seymour Cray); editorial board changes. IEEE Computational Science & Engineering, 4(1):1, 4, 6, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c1001.pdf.

Taflove:1995:REE

[Taf95] Allen Taflove. Reinventing electromagnetics: Emerging applications for FD-TD computation. IEEE Computational Science & Engineering, 2(4): 24–34, Winter 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c4024abs.htm.

Taylor:1994:SRS

[Tay94] Peter Taylor. Site report: The San Diego Supercomputer Center. IEEE Computational Science & Engineering, 1(3):10–14, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Taylor:1996:IVR

[TCD^{+}96] Valerie E. Taylor, Jian Chen, Terrence L. Disz, Michael E. Papka, and Rick Stevens. Interactive virtual reality in simulations: Exploring lag time. IEEE Computational Science & Engineering, 3(4):46–54, Winter 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

                                        URL http://www.computer.
                                        org/cse/cs1998/c4046abs.htm.

Tezduyar:1994:SRA

[Tez94] Tayfun E. Tezduyar. Site report: The Army High Performance Computing Research Center. IEEE Computational Science & Engineering, 1(2):6–8, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Truhlar:1994:CWW

[Tru94a] Donald G. Truhlar. Conferences & workshops: Watching the molecules move. IEEE Computational Science & Engineering, 1 (4):76–78, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Truhlar:1994:EPC

[Tru94b] Donald G. Truhlar. Editors’ perspectives: Computational chemistry: Many roads to travel. IEEE Computational Science & Engineering, 1(3):80–81, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Truhlar:1997:CWC

[Tru97a] Donald G. Truhlar. Conferences & workshops: Computational Drug-Design Workshop. IEEE Computational Science & Engineering, 4(4):84–87, October/December 1997. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1997/pdf/c4084.
pdf.

Truhlar:1997:CWI

[Tru97b] Donald G. Truhlar. Conferences & workshops: International Conference on Parallel Computing. IEEE Computational Science & Engineering, 4(1):98–100, January/March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1098.pdf.

Truhlar:1998:CWQ

[Tru98] Donald G. Truhlar. Conferences & workshops: Quantum Catalysis. IEEE Computational Science & Engineering, 5(2):98–100, April/June 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c2097.pdf.

Umar:1994:EPC

[USO94] Verena Meiser Umar, Michael R. Strayer, and Carl Edward Oliver. Educational programs in CSE: An on-line introduction to computational science via CSEP. IEEE Computational Science & Engineering, 1(4):6–9, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

VandeVelde:1995:BNR

[Van95] Erik F. Van de Velde. Book news & reviews: Studies in Computational Science: Parallel Programming Paradigms by Per Brinch Hansen. IEEE Computational Science & Engineering, 2(4):85– 87, Winter 1995. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Volakis:1995:ECM

[VK95] John L. Volakis and Leo C. Kempel. Electromagnetics: Computational methods and considerations. IEEE Computational Science & Engineering, 2 (1):42–57, Spring 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c1042abs.htm.

Voigt:1994:FCH

[Voi94a] Robert G. Voigt. Focus on CSE:

HPCC/IITA /NII: The alphabet soup of US high-performance computing. IEEE Computational Science & Engineering, 1(2):80, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Voigt:1994:FCI

[Voi94b] Robert G. Voigt. Focus on CSE:

IITA: Access ramp to the information highway. IEEE Computational Science & Engineering, 1 (3):91, Fall 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Voigt:1994:FCN

[Voi94c] Robert G. Voigt. Focus on CSE: NSF plans response to blueribbon panel. IEEE Computational Science & Engineering, 1 (1):84, Spring 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Volakis:1996:CWA

[Vol96] John L. Volakis. Conferences & workshops: 1996 ACES Conference. IEEE Computational Science & Engineering, 3(3):83–84, Fall 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Vrahatis:1997:BNR

[Vra97] Michael N. Vrahatis. Book news & reviews: Ordinary Differential Equations in Theory and Practice: by R. M. M. Matheij and J. Molenaar. IEEE Computational Science & Engineering, 4(4):73–74, October/December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c4073.pdf.

vanEngelen:1997:TWF

[vWC97] Robert van Engelen, Lex Wolters, and Gerard Cats. Tomorrow’s weather forecast: Automatic code generation for atmospheric modeling. IEEE Computational Science & Engineering, 4(3):22–31, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c3022.pdf;   http:
//www.computer.org/cse/cs1998/
c3022abs.htm.

Wahler:1996:CES

[Wah96] Marni Wahler. CSE education: Students get hands-on research experience at SDSC. IEEE Computational Science & Engineering, 3(3):13–16, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Warren:1995:HDM

[War95] James A. Warren. How does a metal freeze? A phasefield model of alloy solidification. IEEE Computational Science & Engineering, 2(2): 38–49, Summer 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c2038abs.htm.

Webb:1997:LER

[Web97] Peter Webb. Letter to the Editor: Response to Wilson: Teach science and software engineering with Matlab. IEEE Computational Science & Engineering, 4(2):4–5, April/June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c2004.pdf.

Weerawarana:1996:BNR

[Wee96]  Sanjiva Weerawarana.  Book

news & reviews: Mathematica as a Tool by Stephan Kaufmann; Mathematica for Scientists and Engineers by Thomas B. Bahder. IEEE Computational Science & Engineering, 3(2):93, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Westropp:1996:SRP

[Wes96] John C. Westropp. Site report:

Pittsburgh Supercomputing Center. IEEE Computational Science & Engineering, 3(1):8–12, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Wijshoff:1996:TSL

[Wij96] Harry Wijshoff. Taking stock, looking ahead: Part 2. programming without bothering about data structures? IEEE Computational Science & Engineering, 3(3):67–68, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (elec-

tronic). URL http://hp137a.
wi.leidenuniv.nl:8080/blas-
service/blas.html;   http:/
/www.computer.org/cse/cs1998/
c3067abs.htm.

Wilson:1994:QCR

[Wil94]  Kenneth G. Wilson.  From

quantum chromodynamics to redesigning education. IEEE Computational Science & Engineering, 1(4):16–18, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Wilson:1996:WSCa

[Wil96a] Gregory V. Wilson. What should computer scientists teach to physical scientists and engineers? IEEE Computational Science & Engineering, 3 (2):46–54, Summer 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c2046abs.htm.

Wilson:1996:WSCb

[Wil96b] Gregory V. Wilson. What should computer scientists teach to physical scientists and engineers? part 2. reply. IEEE Computational Science & Engineering, 3(3):62, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3062abs.htm.

Woodward:1994:ACS

[Woo94] Paul R. Woodward. About the cover: Superfine grids for turbulent flows. IEEE Computational Science & Engineering, 1 (4):4–5, Winter 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Winstead:1995:PCE

[WPM95] Carl Winstead, Howard Pritchard, and Vincent McKoy. Parallel computation of electronmolecule collisions. IEEE Computational Science & Engineering, 2(3):34–42, Fall 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3034abs.htm.

Wolfson:1997:GHO

[WR97] Haim J. Wolfson and Isidore Rigoutsos. Geometric hashing: An overview. IEEE Computational Science & Engineering, 4(4):10–21, October/December 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c4010.pdf;   http:
//www.computer.org/cse/cs1998/
c4010abs.htm.

Wright:1995:EPM

[Wri95] Margaret H. Wright. Editor’s perspective: Medical Applications and Optimization. IEEE Computational Science & Engineering, 2(4):96, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Xu:1997:BNR

[Xu97] Jinchao Xu. Book news & reviews: Domain Decomposition by B. Smith, P. Björstad, and W. Gropp. IEEE Computational Science & Engineering, 4(3):85, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1997/pdf/c3083.pdf.

Zachary:1998:CEI

[Zac98] Joseph L. Zachary. CSE education: An introduction to scientific programming. IEEE Computational Science & Engineering, 5(4):6–10, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

                                        dlib.computer.org/cs/books/
                                        cs1998/pdf/c4006.pdf.

Zhang:1997:DED

[ZDD97] Xiaodong Zhang, Sandra G. Dykes, and Hong Deng. Distributed edge detection: Issues and implementations. IEEE Computational Science & Engineering, 4(1):72–82, January/ March 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c1072.pdf;   http:
//www.computer.org/cse/cs1998/
c1072abs.htm.

Zlatev:1994:MLR

[ZDG94] Zahari Zlatev, Ivan Dimov, and Krassimir Georgiev. Modeling the long-range transport of air pollutants. IEEE Computational Science & Engineering, 1(3):45– 52, Fall 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Zhou:1995:SDS

[ZF95] Jing-Rong Zhou and David K.

Ferry. 3D simulation of deepsubmicron devices: How impurity atoms affect conductance. IEEE Computational Science & Engineering, 2(2): 30–37, Summer 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c2030abs.htm.

Zhu:1995:ACE

[Zhu95] Ting-Ting Zhu. About the cover:

Explosive results: Simulating a four-stroke engine cycle. IEEE Computational Science & Engineering, 2(3):4–5, Fall 1995. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Zima:1996:TSL

[Zim96] Hans P. Zima. Taking stock, looking ahead: Part 2. highperformance languages for parallel computing. IEEE Computational Science & Engineering, 3(3):63–65, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://www.computer.
org/cse/cs1998/c3063abs.htm.

Zhang:1998:MHP

[ZKJ^{+}98] Zengyan Zhang, Satya N. V. Kalluri, Joseph JáJá, Shunlin Liang, and John R. G. Townshend. Models and highperformance algorithms for global BRDF retrieval. IEEE Computational Science & Engineering, 5(4):16–29, October/December 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1998/pdf/c4016.pdf;   http:
//www.computer.org/cse/cs1998/
c4016abs.htm.

Zhu:1998:VRC

[ZM98] Shao-Min Zhu and Phillip John McKerrow. Visualizing reflection off a curved surface. IEEE Computational Science & Engineering, 5(4):30–39, October/December 1998. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

URL http://dlib.computer.
org/cs/books/cs1998/pdf/c4030.
pdf;   http://www.computer.
org/cse/cs1998/c4030abs.htm.

Zeigler:1997:DEH

[ZMKB97] Bernard P. Zeigler, Yoonkeon Moon, Doohwan Kim, and George Ball. The DEVS environment for high-performance modeling and simulation. IEEE Computational Science & Engineering, 4(3):61–71, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://

dlib.computer.org/cs/books/
cs1997/pdf/c3061.pdf;   http:
//www.computer.org/cse/cs1998/
c3061abs.htm.